### v1

# ComSD: Balancing Behavioral Quality and Diversity in Unsupervised Skill Discovery

Abstract
Learning diverse and qualified behaviors for utilization and adaptation without supervision is a key ability of intelligent creatures. Ideal unsupervised skill discovery methods are able to produce diverse and qualified skills in the absence of extrinsic reward, while the discovered skill set can efficiently adapt to downstream tasks in various ways. Maximizing the Mutual Information (MI) between skills and visited states can achieve ideal skill-conditioned behavior distillation in theory. However, it’s difficult for recent advanced methods to well balance behavioral quality (exploration) and diversity (exploitation) in practice, which may be attributed to the unreasonable MI estimation by their rigid intrinsic reward design. In this paper, we propose Contrastive multi-objectives Skill Discovery (ComSD) which tries to mitigate the quality-versus-diversity conflict of discovered behaviors through a more reasonable MI estimation and a dynamically weighted intrinsic reward. ComSD proposes to employ contrastive learning for a more reasonable estimation of skill-conditioned entropy in MI decomposition. In addition, a novel weighting mechanism is proposed to dynamically balance different entropy (in MI decomposition) estimations into a novel multi-objective intrinsic reward, to improve both skill diversity and quality. For challenging robot behavior discovery, ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot. On numerical evaluations, ComSD exhibits state-of-the-art adaptation performance, significantly outperforming recent advanced skill discovery methods across all skill combination tasks and most skill finetuning tasks. Codes will be released at https://github.com/liuxin0824/ComSD.
1 Introduction&Research Background
Reinforcement Learning (RL) has proven its effectiveness in learning useful task-specific skills in the presence of extrinsic rewards  (Mnih et al., 2015; Levine et al., 2016; Ding et al., 2021; Narvekar et al., 2017; Li et al., 2019). The success of unsupervised learning in computer vision (Chen et al., 2020; Caron et al., 2020) and natural language processing (Brown et al., 2020; Devlin et al., 2018) further benefits task-specific RL with complex input (Laskin et al., 2020; Kostrikov et al., 2020; Yarats et al., 2021a; Stooke et al., 2021) by improving representation learning. However, the agents trained with lots of efforts are always hard to generalize their knowledge to novel tasks due to the task-specific supervision of extrinsic reward (Stooke et al., 2021). In addition to the generalization, intelligent agents should also be able to explore environments and learn different useful behaviors without any extrinsic supervision, like human beings. For the reasons above, unsupervised skill discovery is proposed and becomes a novel research hotspot (Gregor et al., 2016; Eysenbach et al., 2018). As a branch of unsupervised RL  (Pathak et al., 2017; 2019; Burda et al., 2018), unsupervised skill discovery  (Park et al., 2021; 2023; Strouse et al., 2021) also designs task-agnostic rewards and achieves unsupervised pre-training with these rewards, which guarantees the task-agnostic downstream generalization. The main difference is that skill discovery requires extra input as conditions, named skill vectors. They aim to discover useful task-agnostic policies that are distinguishable by skill vectors. The discovered skills can be implemented for downstream tasks in various ways, including skill finetuning, skill combination, skill imitation, and so on.
Currently, one of the most popular and effective classes of skill discovery is based on Mutual Information (MI) maximization (Eysenbach et al., 2018; Sharma et al., 2019; Liu & Abbeel, 2021a; Yang et al., 2023). Most of them try to design intrinsic rewards to estimate and optimize the MI between skill vectors and visited states. With RL maximizing intrinsic reward expectation, correlations between skills and states are distilled, i.e., different useful skills are discovered. Recent MI-based methods exhibit considerable results in different fields, including map exploration (Campos et al., 2020), robotic manipulation (Plappert et al., 2018) and video games (Bellemare et al., 2013). However, when facing challenging robot locomotion where the space of state and action are continuous, complex, and non-linear, existing advanced methods can’t balance the behavioral quality (exploration) and diversity (exploitation) well. Specifically, they can either only learn different static postures of lazy exploration (Eysenbach et al., 2018; Liu & Abbeel, 2021a) or only produce highly dynamic behaviors that are homogeneous and indistinguishable (i.e., insufficient exploitation) (Laskin et al., 2022b), while an ideal robot behavior set should contain diverse behaviors at different activity levels, enabling efficient downstream adaptation in various ways. Some recent works have also noticed this issue (Yang et al., 2023), as we have. However, they only address it on 2D exploration problems but remain powerless for the difficult multi-joint robot behavior discovery.
Figure 1: ComSD discovers different useful unsupervised skills through a novel multi-objective intrinsic reward. Skill combination and skill finetuning together provide comprehensive skill evaluation, considering both behavioral quality and diversity. Our ComSD exhibits state-of-the-art adaptation performance on both kinds of downstream tasks, which recent advanced methods cannot.
We attribute the above quality-versus-diversity conflict to their unreasonable MI estimation and their rigid intrinsic reward design. In this paper, we propose ComSD, which tries to mitigate this conflict by a more reasonable MI estimator and a dynamic weighting algorithm for intrinsic reward design. Concretely, ComSD decomposes the MI between state transitions and skill vectors into the negative skill-conditioned entropy and state entropy, designing a novel intrinsic reward for MI estimation and optimization (with RL). For conditioned entropy, ComSD proposes to employ the contrastive learning result between skills and corresponding state transitions as a better estimation. For state entropy, ComSD follows recent works (Liu & Abbeel, 2021a; Yarats et al., 2021b; Laskin et al., 2022b), choosing a popular particle-based estimation (Liu & Abbeel, 2021b). Moreover, a novel Skill-based Multi-objective Weighting (SMW) mechanism is proposed to dynamically balance the above two entropy estimations into a novel multi-objective intrinsic reward, to encourage both exploration and exploitation. ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels for challenging multi-joint robots, which recent advanced methods cannot.
For comprehensive numerical evaluation, two different downstream adaptation tasks are employed: skill combination and skill finetuning. Skill finetuning is widely used recently (Laskin et al., 2022b; Yuan et al., 2022; Yang et al., 2023) on URLB (Laskin et al., 2021), where a target skill is chosen from the skill set and further finetuned with extrinsic reward. However, it’s one-sided to judge the whole skill set with only one skill, and the process of finetuning also introduces much uncertainty. In addition, behavioral diversity is also ignored in skill finetuning. To this end, we further employ another skill combination (Eysenbach et al., 2018) to evaluate both behavioral quality and diversity, where the learned skills are frozen and a meta-controller is trained to combine the learned skills for downstream task achievement. ComSD outperforms all baselines on skill combination significantly and is competitive with state-of-the-art methods on skill finetuning, as shown in Figure 1.
Our contributions can be summarized as follows: (i) We propose ComSD, a novel unsupervised skill discovery algorithm that discovers different useful behaviors by maximizing the MI objective between skill vectors and state transitions. (ii) ComSD proposes to employ the contrastive learning result between skills and states as a more reasonable conditioned entropy estimation in MI decomposition. (iii) A novel weighting algorithm, SMW, is presented to produce a novel multi-objective intrinsic reward that improves skill diversity and quality simultaneously. (iv) Comprehensive numerical evaluation and detailed analysis show that ComSD outperforms all recent advanced methods, exhibiting state-of-the-art downstream adaptation ability. It can produce a qualified skill set consisting of diverse behaviors for challenging multi-joint robots, which recent advanced methods cannot.
2 Preliminaries
2.1 Problem Definition
Unsupervised skill discovery algorithms aim at discovering a useful and diverse set of agent behaviors in the absence of extrinsic reward, i.e., training a skill-conditioned policy in a task-agnostic environment. Concretely, a reward-free Markov Decision Process (MDP) (Bellman, 1957) is considered and defined as $\mathcal{M}^{free}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the distribution of the next state given the current state and action, $\gamma$ is the discount factor, and $d_{0}$ is the distribution of the initial state. What skill discovery algorithms do is to define an intrinsic reward $r^{intr}$ and augment $\mathcal{M}^{free}$ to a intrinsic-reward MDP $\mathcal{M}^{intr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{intr},\gamma,d_{0})$. With $z\sim p(z)$ given by discovery algorithms, the skill-conditioned policy $\pi(a|s,z)$ can be obtained by RL over the $\mathcal{M}^{intr}$. As an example, we provide the pseudo-code of our ComSD in Appendix A.
To evaluate the adaption ability of discovered skills (i.e., the skill-conditioned agent $\pi(a|s,z)$), two adaptation evaluations: skill combination and skill finetuning are employed on each task-specific downstream task. A downstream task can be described as an extrinsic-reward MDP $\mathcal{M}^{extr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{extr},\gamma,d_{0})$, where $r^{extr}$ denotes the task-specific extrinsic reward. In skill combination, the learned $\pi(a|s,z)$ is frozen, and an meta-controller $\pi^{\prime}(z|s)$ is required to choose and combine skill vectors for $\pi(a|s,z)$ automatically on downstream tasks. Concretely, over the original task $\mathcal{M}^{extr}$, skill combination can be defined as $\mathcal{M}^{extr^{\prime}}=(\mathcal{S},\mathcal{Z},\mathcal{P^{\prime}},r^{extr^{\prime}},\gamma,d_{0})$, which regards different skill vectors $z\sim\mathcal{Z}$ as its actions and correspondingly changes its transition model and reward function. The meta-controller $\pi^{\prime}(z|s)$ is trained over $\mathcal{M}^{extr^{\prime}}$ by RL and then used for evaluation of the learned $\pi(a|s,z)$. In skill finetuning, a target skill vector $z_{i}$ is chosen. The corresponding policy $\pi(a|s,z_{i})$ is further finetuned on the $\mathcal{M}^{intr}$ for a few steps and serves as the evaluation of the discovered $\pi(a|s,z)$. The pseudo-codes of two adaptation tasks are provided in Appendix A.
2.2 Mutual Information Objective
As most skill discovery algorithms do, our ComSD also tries to maximize the Mutual Information (MI) objective between states $\tau(s)$ and skills $z$. In general, $\tau(s)$ can be (i) maintaining the original state $\tau(s)=s$, (ii) concatenating the neighboring state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$, or (iii) using the whole trajectory $\tau(s)={\rm concat}(s^{1},...,s^{t})$. Following the recent advanced method (Laskin et al., 2022b), we define $\tau$ as state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$ throughout this paper. The MI objective $I(\tau;z)$ can be decomposed into the form of Shannon entropy in two ways:
$I(\tau;z)=-H(z|\tau)+H(z)$ (1)
$I(\tau;z)=-H(\tau|z)+H(\tau)$ (2)
where the $I(\cdot;\cdot)$ denotes the MI function and $H(\cdot)$ denotes the Shannon entropy throughout the paper. Most classical MI-based algorithms (Gregor et al., 2016; Achiam et al., 2018; Eysenbach et al., 2018; Lee et al., 2019)are based on the first decomposition Eq. 1. A uniform random policy can guarantee the maximum of the skill entropy $H(z)$ while the trainable discriminators estimating the conditioned entropy $H(z|\tau)$ are employed to calculate the intrinsic reward for unsupervised RL. However,  Laskin et al. (2021; 2022b) have found it hard for these methods to guarantee robot behavioral activity and exploration without external balancing signals. To this end, some recent works (Sharma et al., 2019; Liu & Abbeel, 2021a; Laskin et al., 2022b) focus on the second decomposition Eq. 2. They explicitly optimize the state entropy $H(\tau)$ by exploratory intrinsic rewards, achieving considerable performance on several tasks, including robotic manipulation, video games, and robot locomotion. Following these, we choose Eq. 2 as our optimization target.
3 Unsupervised Skill Discovery by ComSD
Figure 2: Left: The multi-objective intrinsic reward design of ComSD. Right: In SMW, $\beta$ is linear related to $\rm flag\it(z)$ with different slopes in different region.
In this section, we detail how ComSD designs a novel multi-objective intrinsic reward based on Eq. 2. In Section 3.1, we propose and describe how to employ contrastive learning to design an exploitation reward for the negative conditioned entropy estimation. In Section 3.2, we illustrate the way to estimate the state entropy by a particle-based exploration reward, which follows the recent advanced methods. In Section 3.3, we describe why and how to balance the above two sub-intrinsic rewards into a multi-objective intrinsic reward with proposed SMW. The overall intrinsic reward design of ComSD is visualized in Figure 2.
3.1 Conditioned Entropy Estimation via Contrastive Learning
In decomposition Eq. 2, increasing the negative conditioned entropy $-H(\tau|z)$ intuitively encourages the agent to access familiar and visited state pairs according to the corresponding skills. This helps the agent exploit skill-specific information, i.e., increase the diversity of learned skills. Due to the unavailable conditioned distribution, we can’t directly maximize the $-H(\tau|z)$ but try to maximize its lower bound. First, $-H(\tau|z)$ can be decomposed as:
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau|z)=\sum_{\tau,z}p(\tau,z)\log\frac{p(\tau,z)}{p(z)}$ (3)
$=\sum_{\tau,z}p(\tau,z)(\log p(\tau,z)-\log p(z)),$
where $p(\cdot)$ denotes probability throughout the paper. In practice, we use a uniform distribution to generate the skill vectors. Therefore, the term $\rm log\it p(z)$ becomes a constant, and we denote it as $c$:
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)(\rm log\it p(\tau,z)-c)=\sum_{\tau,z}p(\tau,z)\rm log\it p(\tau,z)-\sum_{\tau,z}p(\tau,z)\cdot c$ (4)
$=\sum_{\tau,z}p(\tau,z)logp(\tau,z)-c\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z).$
Then we can obtain the variational lower bound:
$-H(\tau|z)$ $\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z)$ (5)
$=\sum_{\tau,z}p(\tau,z)\log p(\tau,z)+\sum_{\tau,z}p(\tau,z)\log q(\tau,z)-\sum_{\tau,z}p(\tau,z)\log q(\tau,z)$
$=\sum_{\tau,z}p(\tau,z)\log q(\tau,z)+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$
$=\mathbb{E}_{\tau,z}[\log q(\tau,z)]+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))\geq\mathbb{E}_{\tau,z}[\log q(\tau,z)],$
where $q(\tau,z)$ is employed to estimate the unavailable $p(\tau,z)$. To optimize $-H(\tau|z)$, we need to minimize the KL divergence $D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$ and maximize the expectation $\mathbb{E}_{\tau,z}[\log q(\tau,z)]$.
For KL divergence minimization, the distribution $q$ should be as positively correlated with the probability function $p$ as possible. We define $q$ as the exponential inner product between the state embeddings and skill embeddings:
$q(\tau,z)=\exp\frac{f_{\theta_{1}}(\tau)^{T}\cdot f_{\theta_{2}}(z)}{||f_{\theta_{1}}(\tau)^{T}||\cdot||f_{\theta_{2}}(z)||T},$ (6)
where $f_{\theta_{1}}(\cdot)$ is the state encoder, $f_{\theta_{2}}(\cdot)$ is the skill encoder, T is the temperature, and $\exp(\cdot)$ makes $q$ non-negative like $p$. The above two encoders are updated by gradient descent after computing the NCE loss (Gutmann & Hyvärinen, 2010) in contrastive learning (Chen et al., 2020):
$\mathcal{L}_{NCE}=\log q(\tau_{i},z_{i})-\log\frac{1}{N}\sum_{j=1}^{N}q(\tau_{j},z_{i}),$ (7)
where $\tau_{i}$ is sampled by skill $z_{i}$ and $\tau_{j}$ is sampled by other skills. In most circumstances, the probability $p(\tau_{i},z_{i})$ is much larger than $p(\tau_{j},z_{i})$. Maximizing $\mathcal{L}_{NCE}$ increases the value of $q(\tau_{i},z_{i})$ and decreases the value of $q(\tau_{j},z_{i})$, which actually shrinks the difference between $p$ and $q$.
For expectation maximization, we regard the optimization objective as a part of our unsupervised intrinsic reward, employing RL to optimize it:
$r^{intr}_{exploitation}\propto q(\tau_{i},z_{i}).$ (8)
3.2 Particle-based State Entropy Estimation
In decomposition Eq. 2, increasing the state entropy $H(\tau)$ encourages the agent to explore more widely and visit more state transitions. Following  Liu & Abbeel (2021a); Yarats et al. (2021b); Laskin et al. (2022b), we employ a particle-based entropy estimation proposed by  Liu & Abbeel (2021b). Concretely, $H(\tau)$ can be estimated by the Euclidean distance between each particle ($\tau$ in our paper) and its all $k$-nearest neighbor (Singh et al., 2003) in the latent space:
$H(\tau)\approx H_{particle}(\tau)\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{buffer}^{k-nn}(h_{i})}\log||h_{i}-h_{i}^{nn}||,$ (9)
where $h_{i}$ can be any forms of encoded $\tau_{i}$ and $N_{buffer}^{k-nn}(h_{i})$ denotes neighbor set. Following  Laskin et al. (2022b), we choose the state encoder in contrastive representation learning and calculate this entropy over sampled RL training batch. The exploration reward is defined as the following:
$r^{intr}_{exploration}\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{batch}^{k-nn}(h_{i})}\log||h_{i}-h_{i}^{nn}||,\ \ where\ \ h_{i}=f_{\theta_{1}}(\tau_{i}).$ (10)
3.3 Skill-based Multi-objectives weighting
With the negative conditioned entropy $-H(\tau|z)$ estimated by $r^{intr}_{exploitation}$ and the state entropy estimated by $r^{intr}_{exploration}$, a naive way for intrinsic reward design is to employ a fixed coefficient $\alpha$ to scale the two terms ($r^{intr}_{naive}=r^{intr}_{exploration}+\alpha r^{intr}_{exploitation}$). However, we find it impossible to choose a proper fixed coefficient to balance them well especially in challenging multi-joint robot behavior discovery, which is because of a severe quality-versus-diversity conflict between the above two sub-intrinsic rewards. Specifically, when employing $r^{intr}_{exploration}$ alone, the agents can only learn dynamic behaviors of high activity but not easing movements or static postures, and the learned skills are highly homogeneous and indistinguishable by skill vectors (i.e., unsatisfactory exploitation). The naive intervention of $r^{intr}_{exploitation}$ can significantly improve the behavioral diversity, but it will simultaneously cause the lazy exploration, making the robots unable to discover active behaviors.
We propose Skill-based Multi-objectives Weighting (SMW), a simple but effective dynamic weighting mechanism for the issue above. SMW aims to take advantage of the two sub-rewards by setting different optimization objectives for different skill vectors. Concretely, it designs another changeable coefficient $\beta$ which is dynamically adjusted according to skill vectors. $\beta(z)$ is defined as:
[…]
The $\rm flag\it(\cdot)$ can be any function to map the high-dimensional skill vectors into 1-dimensional flag variables. We find that simply employing the first dimension of skill vectors (i.e., $\rm flag\it(z)=I_{1}^{T}\cdot z,\ {\rm where}\ I_{1}=(1,0,...,0)$) can perform well. The fixed hyper-parameters $w_{high},w_{low},f_{high},f_{low}$ and the clamp function $\rm clamp\it(\cdot)$ make the dynamic coefficient $\beta$ linearly related to flag variables $\rm flag\it(z)$ with different slopes in different regions. We clearly visualize the relation between $\beta$ and $\rm flag\it(z)$ in Figure 2. With SMW, the multi-objective intrinsic reward of ComSD is defined as:
$r^{intr}_{ComSD}=r^{intr}_{exploration}+\alpha\cdot\beta(z)\cdot r^{intr}_{exploitation}.$ (12)
Over the intrinsic-reward MDP $\mathcal{M}^{intr}$, the skill-conditioned policy $\pi(a|s,z)$ can be trained by RL with $z\sim p(z)$. Algorithm 1 in Appendix A provides the full pseudo-code of ComSD.
4 Skill Evaluation and Analysis
4.1 Experimental Setup
Environments. On the importance of environments: OpenAI Gym (Brockman et al., 2016) and DMControl (Tassa et al., 2018) are two most popular continuous robot locomotion benchmarks. In Gym, the episode is ended when agents lose their balance, while the episode length in DMControl is fixed. Laskin et al. (2022b) found this difference makes DMControl much harder for reward-free exploration since agents have to learn balance by themselves without any external signals.
Following recent advanced methods (Laskin et al., 2022b; Yarats et al., 2021b; Liu & Abbeel, 2021a; b; Liu et al., 2023), we employ 16 downstream tasks of 4 domains from URLB (Laskin et al., 2021) and DMControl (Tassa et al., 2018) for skill evaluation. The domains Walker, Quadruped, Cheetah and Hopper are recognized as the most representative and challenging multi-joint robot environments, each of which contains 4 totally diverse and challenging downstream locomotion tasks.
Evaluations. In each domain, all the methods, including ComSD pre-train their agents for 2M environment steps with their respective intrinsic rewards. After unsupervised pre-training, the behavioral diversity and quality will be fully and reasonably evaluated on two adaptation tasks: skill combination (Eysenbach et al., 2018) and skill combination in URLB (Laskin et al., 2021) respectively, across all 16 downstream tasks with extrinsic reward. It means a total of 32 numerical results are employed for one method’s evaluation. DDPG (Lillicrap et al., 2015) is chosen as the backbone RL algorithm for all methods throughout the paper. The detailed settings of the skill combination and skill finetuning are provided along with the experimental results in Section 4.2.
Baselines. We compare our ComSD with five recent advanced skill discovery algorithms that are popular for robot behavior discovery. They are BeCL (Yang et al., 2023), CIC (Laskin et al., 2022b), APS (Liu & Abbeel, 2021a), SMM (Lee et al., 2019), and DIAYN (Eysenbach et al., 2018). All the methods try to optimize their MI target by designing intrinsic rewards for unsupervised RL. In addition to BeCL, they all try to maximize $I(\tau,z)$, where DIAYN and SMM choose the decomposition Eq. 1 while APS, CIC, and our ComSD choose decomposition Eq. 2. BeCL proposes a novel MI objective $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill. We provide a detailed description of these baselines in Appendix B.
The close baselines to our ComSD are CIC and APS. Apart from SMW, a novel weighting mechanism, ComSD still differs significantly from these methods. For CIC, ComSD follows it in state entropy estimation but first proposes to employ contrastive results for explicit state entropy maximization. APS and ComSD both explicitly optimize the state entropy and conditioned entropy, while ComSD utilizes contrastive learning for a more reasonable estimation. In addition, BeCL also employs contrastive learning like ComSD, but their contrastive components are different.
4.2 Main Results
Figure 3: The training curve of all 6 methods on 16 downstream skill combination tasks. ComSD outperforms all baselines significantly across all 16 tasks, demonstrating that ComSD discovers much more diverse and qualified behaviors than other methods for challenging multi-joint robots.
Skill Combination. In skill combination, we train another meta-controller $\pi^{\prime}(z|s)$ which selects the discovered skills automatically to achieve downstream tasks. The discovered skill-conditioned agent $\pi(a|s,z)$ is frozen, and $\pi^{\prime}(z|s)$ is trained by RL with extrinsic reward (see Section 2.1 and Appendix A.2 for a detailed problem definition). This adaptation task can effectively evaluate both the diversity and quality of the discovered skill set. For each method, 2M environment steps RL are allowed to train their meta-controllers. All the hyper-parameters of meta-controller and RL settings are shared across all the methods for a fair comparison, which we detail in Appendix C. The results of ComSD are obtained over 6 different random seeds.
The training curves are shown in Figure 3, which demonstrates that ComSD significantly outperforms all five baselines across all 16 downstream tasks. In skill combination, the better adaptation performance means that the behaviors discovered by ComSD are of higher diversity and quality than other methods. CIC and APS are the two most competitive baselines. Both of them and our ComSD employ the decomposition Eq. 2, which indicates the importance of explicitly maximizing state entropy for discovering robot behaviors. An interesting phenomenon is that CIC always gets a higher initial score than APS and ComSD but has no upward trend, which coincides with the fact that CIC can only find homogeneous behaviors at a high activity level. By contrast, APS and ComSD explicitly maximize the conditioned entropy through intrinsic rewards, greatly improving the diversity of learned skill sets and achieving higher final scores than CIC. Compared with APS, ComSD performs better due to the following two reasons: First, ComSD employs contrastive learning for a better estimation of both state entropy and conditioned entropy. Second, the proposed SMW enables ComSD to learn qualified behaviors at different activity levels, thus gaining better adaptation ability.
Skill Finetuning. In skill finetuning, a skill vector $z_{i}$ is chosen, and the corresponding skill $\pi(\cdot|s,z_{i})$ is finetuned with the extrinsic reward (see Section 2.1 and Appendix A.3 for a detailed problem definition). 100k environment steps with extrinsic reward are allowed for each method, where a proper skill vector is selected in the first 4k steps and the chosen skill is finetuned in another 96k steps. For skill selection, previous works employ random sampling (Yang et al., 2023), fixed choice (Laskin et al., 2022b; a), or reward-based choice (Liu & Abbeel, 2021a). We follow the official implementation of  Laskin et al. (2022b), employing a fixed mean skill with the first dimension set to 00. All the hyper-parameters of neural architecture and RL are shared across different methods for a fair comparison. We show the detailed experimental settings in Appendix C. The results of ComSD are obtained over 10 different random seeds.
Table 1: The numerical results of all 6 methods on 16 skill finetuning downstream tasks. ComSD exhibits competitive adaptation performance compared with state-of-the-art methods (CIC and BeCL). In skill finetuning, only one skill is used for evaluation. The behavioral diversity is also ignored. It is one-sided to employ skill finetuning alone for evaluation, like recent works.
Domain Task DIAYN SMM APS CIC BeCL ComSD(ours)
Walker run 242±11 430±26 257±27 450±19 387±22 447±64
stand 860±26 877±34 835±54 959±2 952±2 962±9
flip 381±17 505±26 461±24 631±34 611±18 630±41
walk 661±26 821±36 711±68 885±28 883±34 918±32
Quadruped run 415±28 220±37 465±37 445±36 482±105 500±103
stand 706±48 367±42 714±50 700±55 789±142 824±86
walk 406±64 184±26 602±86 621±69 707±197 735±140
jump 578±46 298±39 538±42 565±44 610±134 686±66
Hopper hop 3±4 5±7 1±1 59±60 5±7 40±35
flip 7±8 29±16 3±4 96±64 13±15 61±47
hop backward 9±28 29±57 2±0 172±64 40±72 92±105
flip backward 2±1 19±34 10±23 154±70 22±36 59±63
Cheetah run 151±72 506±35 381±41 483±32 380±35 432±37
flip 615±78 711±6 648±82 730±13 701±30 660±52
run backward 368±15 473±19 392±41 452±11 400±20 458±9
flip backward 477±108 679±7 518±103 678±93 643±102 685±55
The numerical results on 16 downstream tasks are shown in Table 1. CIC and BeCL are the state-of-the-art methods for skill finetuning. ComSD performs comparable or better on 10/16 tasks than CIC and outperforms BeCL across 15/16 tasks, also achieving state-of-the-art adaptation performance. Actually, the score of the initial few steps has little to do with the final score in state-based locomotion RL, which means it’s hard to find a proper skill from a large skill set within only 4k steps. This explains why CIC and BeCL use random or fixed choices. Furthermore, higher behavioral diversity causes higher choice difficulty in ComSD. ComSD discovers the most diverse behaviors (state-of-the-art adaptation on skill combination) while simultaneously achieving competitive performance on skill finetuning with a fixed skill, which further shows its superior behavioral quality.
Figure 4: Numerical ablation experiments on (left) two skill combination tasks and (right) two skill finetuning tasks. Contrastive estimation of negative conditioned entropy and SMW are both necessary for ComSD to achieve the most competitive results on both kinds of adaptation tasks.
Figure 5: Skill activity analysis of our ComSD, ComSD w/o SMW, and all baselines. We estimate the skill activity by visited state entropy (KTD and AKD) and report the activity variance and range of uniformly selected walker skills for each method. ComSD exhibits much higher entropy variance and range, demonstrating that ComSD indeed discovers skills at different activity levels.
4.3 Numerical Ablation Study
In this section, we conduct numerical ablation experiments to show the effectiveness of both contrastive conditioned entropy estimation and SMW. The results are shown in Figure 4, where CIC is actually ComSD w/o explicit conditioned entropy estimation&SMW, and APS is actually ComSD w/o contrastive conditioned entropy estimation&SMW. Compared with implicit maximization (CIC), explicitly maximizing the skill-conditioned entropy (APS and ComSD w/o SMW) improves the training curve slope on skill combination, which means better exploitation, i.e., behavioral diversity. Moreover, the contrastive result is a better conditioned entropy estimation, improving APS on both kinds of adaptation tasks. However, simply maximizing the conditioned entropy explicitly leads to a trivial solution in practice, where the agent gives up continuous movement and exploration. They can only learn static postures but not dynamic behaviors, exhibiting performance drops on skill finetuning compared with CIC. This phenomenon is also observed by Laskin et al. (2022b). SMW effectively alleviates this exploration hurt by setting multiple optimization objectives. It improves behavioral exploitation while maintaining the tendency toward exploration, making ComSD exhibit state-of-the-art performance on both kinds of adaptation tasks. In summary, SMW and contrastive conditioned entropy estimation are both necessary for ComSD.
4.4 Quality Analysis of Learned Skills
Numerical results on 32 downstream tasks have demonstrated the effectiveness and superiority of ComSD. In this section, we directly analyze the learned skills to show that our method indeed discovers behaviors at different activity levels, which previous advanced methods cannot. For each method, we uniformly select eleven walker skills and use the entropy of visited states to represent the skill activity. K-Th-nearest-neighbor Distance (KTD) and All-K-nearest-neighbor Distance (AKD) (Liu & Abbeel, 2021b), are employed as state entropy estimations. KTD computes the k-th-nearest-neighbor Euclidean distance of each state and takes the mean, while AKD considers all the k-nearest neighbors of each state and takes the mean distance. For each method, we report the entropy variance and entropy range of the selected skills in Figure 5. ComSD discovers a skill set with the biggest activity range and variance, demonstrating it indeed produces skills at diverse activity levels, where mainly attributes to multiple optimization targets brought by multi-objective intrinsic reward. We strongly refer readers to Appendix D for additional visualization and analysis.
5 Conclusion
In this paper, we propose a novel unsupervised skill discovery method named ComSD. It improves both behavioral quality and diversity through contrastive learning from a multi-objective perspective. For numerical evaluation, we conduct comprehensive comparisons on different kinds of downstream adaptation tasks, considering behavioral diversity ignored by recent works. The superiority of ComSD over all other baselines is well verified by our employed evaluation. We hope (i) our ComSD can inspire more attention to intrinsic reward weighting algorithms, and (ii) our experimental settings can make a difference for unsupervised skill evaluation in future works.
Appendix
Appendix A Full pseudo-code
A.1 ComSD
We provide the ComSD’s full pseudo-code. It also serves as an example to demonstrate the process of general unsupervised skill discovery.
Algorithm 1 Pseudo-code of ComSD.
###Unsupervised skill discovery by ComSD
Require: Reward-free environment $E_{f}$, the uniform skill distribution $p(z)$, unsupervised pre-training environment steps $I_{p}$, and the RL batch size $I_{b}$.
Initialize: The state encoder $f_{\theta_{1}}(\cdot)$, the skill encoder $f_{\theta_{2}}(\cdot)$, the skill-conditioned policy (actor) $\pi(a|s,z)$, the critic $Q(a,concat(s,z))$, and the replay buffer $D$.
1:  for $t=1,...,I_{p}$ do
2:       Sample a skill vector from uniform distribution $z_{t}\sim p(z)$.
3:       Obtain current action $a_{t}\sim\pi(\cdot|s_{t},z_{t})$ based on current observation $s_{t}$.
4:       Interact with reward-free environment $E_{f}$ with $a_{t}$ to get next observation $s_{t+1}$.
5:       Add the transition $(s_{t},z_{t},a_{t},s_{t+1})$ into replay buffer $D$.
6:       Sample $I_{b}$ transitions from $D$ (after enough data collection).
7:       Compute contrastive learning loss $\mathcal{L}_{NCE}$ shown in Eq. 7 with $f_{\theta_{1}}(\cdot)$ and $f_{\theta_{2}}(\cdot)$.
8:       Use backpropogation to update $f_{\theta_{1}}(\cdot)$ and $f_{\theta_{2}}(\cdot)$.
9:       Compute exploitation reward $r^{intr}_{exploitation}$ with Eq. 8.
10:       Compute exploration reward $r^{intr}_{exploration}$ with Eq. 10.
11:       Compute the final intrinsic reward $r^{intr}_{ComSD}$ with Eq. 12.
12:       Augment sampled transition batch by the $r^{intr}_{ComSD}$.
13:       Use DDPG to update $\pi(a|s,z)$ and $Q(a,concat(s,z))$ over $I_{b}$ intrinsic-reward transitions.
14:  end for
Output: the discovered skills (trained skill-conditioned policy) $\pi(a|s,z)$.
A.2 Skill Combination Evaluation
Algorithm 2 Pseudo-code of skill combination evaluation.
###Adaption evaluation of pre-trained skills by skill combination
Require: Reward-specific environment $E_{s}$, the pre-trained skill-conditioned policy $\pi(a|s,z)$, environment adaption steps $I_{a}$, and the RL batch size $I_{b}$.
Initialize: The meta-controller (actor) $\pi^{\prime}(z|s)$, the critic $Q(z,s)$, and the replay buffer $D$.
1:  Freeze the learned skills $\pi(a|s,z)$.
2:  for $t=1,...,I_{a}$ do
3:       Obtain current skill vector $z_{t}\sim\pi(\cdot|s_{t})$ based on current observation $s_{t}$.
4:       Obtain current action $a_{t}\sim\pi(\cdot|s_{t},z_{t})$ based on $s_{t}$ and $z_{t}$.
5:       Interact with reward-specific environment $E_{s}$ with $a_{t}$ to get next observation $s_{t+1}$ and the extrinsic reward $r_{extr}$.
6:       Add the transition $(s_{t},z_{t},r_{extr},s_{t+1})$ into replay buffer $D$.
7:       Sample $I_{b}$ transition batch from $D$.
8:       Use DDPG to update $\pi^{\prime}(z|s)$ and $Q(z,s)$ over $I_{b}$ transitions.
9:  end for
Output: The performance of $\pi^{\prime}(z|s)$ serves as the skill combination evaluation result.
A.3 Skill Finetuning Evaluation
Algorithm 3 Pseudo-code of skill finetuning evaluation.
###Adaption evaluation of pre-trained skills by skill finetuning
Require: Reward-specific environment $E_{s}$, the pre-trained skill-conditioned policy $\pi(a|s,z)$, environment adaption steps $I_{a}$, skill choice steps $I_{c}$, and the RL batch size $I_{b}$.
Initialize: The critic $Q(a,concat(s,z))$, and the replay buffer $D$.
1:  Choose a skill vector $z_{i}$ in $I_{c}$ steps by your algorithm (e.g., a fixed choice in CIC and ComSD) and save the corresponding $I_{c}$ extrinsic-reward transitions into $D$.
2:  Freeze the chosen skill vector $z_{i}$.
3:  for $t=1,...,I_{a}-I_{c}$ do
4:       Obtain current action $a_{t}\sim\pi(\cdot|s_{t},z_{i})$ based on $s_{t}$ and $z_{i}$.
5:       Interact with reward-specific environment $E_{s}$ with $a_{t}$ to get next observation $s_{t+1}$ and the extrinsic reward $r_{extr}$.
6:       Add the transition $(s_{t},z_{i},a_{t},r_{extr},s_{t+1})$ into replay buffer $D$.
7:       Sample $I_{b}$ transition batch from $D$.
8:       Use DDPG to update $\pi(a|s_{t},z_{i})$ and $Q(a,concat(s,z))$ over $I_{b}$ transitions.
9:  end for
Output: The performance of $\pi(a|s,z_{i})$ serves as the skill finetuning evaluation result.
Appendix B Baseline Details
DIAYN
(Eysenbach et al., 2018) is one of the most classical and original unsupervised skill discovery algorithms, trying to maximize the MI between skills and states. It employs the first MI decomposition, Eq. 1. It uses a discrete uniform prior distribution to guarantee the maximization of skill entropy $H(z)$. The negative state-conditioned entropy $-H(z|s)$ is estimated by a trainable discriminator ${\rm log}p(z|s)$ which computes the intrinsic reward. As a foundational work, it provides several reasonable evaluations of skill adaptation, of which skill finetuning and skill combination are employed in our experiments.
SMM
(Lee et al., 2019) aims to learn a policy for which the state marginal distribution matches a given target state distribution. It optimizes the objective by reducing it to a two-player, zero-sum game between a state density model and a parametric policy. Like DIAYN, it is also based on the first decomposition (Eq. 1) of MI and employs discriminator training. The difference is that SMM explicitly maximizes the state entropy with intrinsic reward, which inspires lots of recent advanced works and our ComSD.
APS
(Liu & Abbeel, 2021a) first employs the second MI decomposition Eq. 2 for a better MI estimation. For state entropy estimation, it employs a popular particle-based entropy estimation proposed by APT (Liu & Abbeel, 2021b), which is proven effective and supports many advanced works (Laskin et al., 2022b; Yarats et al., 2021b) and our ComSD. For skill-conditioned entropy, it chooses the successor feature (Hansen et al., 2019), introducing it into the final intrinsic reward for an explicit maximization. The weight between different entropy estimations is fixed in APS. APS can’t guarantee the behavioral quality (exploration) well. Different from APS, our ComSD employs contrastive learning for better conditioned entropy estimation and designs a novel dynamic weighting algorithm (SMW) to overcome the exploration drop brought by explicit conditioned entropy maximization.
CIC
(Laskin et al., 2022b) is a state-of-the-art robot behavior discovery method. It first introduces contrastive learning (Chen et al., 2020) into unsupervised skill discovery. It chooses the second MI decomposition, Eq. 2 with APT particle-based estimation for state entropy, like APS. The contrastive learning between state transitions and skill vectors is conducted for implicit skill-conditioned entropy maximization. The encoder learned by contrastive learning is further used for APT reward improvement. The behaviors produced by CIC are of high activity but not distinguishable. Different from CIC, we employ the contrastive results as diversity intrinsic rewards for explicit conditioned entropy maximization to improve behavioral diversity, with SMW to balance two entropy estimations for exploratory ability maintenance.
BeCL
(Yang et al., 2023) is another state-of-the-art method on URLB (Laskin et al., 2021) and 2D exploration (Campos et al., 2020). It tries to mitigate the exploitation problem in CIC by a novel MI objective, $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill. It provides theoretical proof to show that their novel MI objective serves as the upper bound of the previous MI objective. However, BeCL can’t generate enough dynamic robot behaviors, and their intrinsic reward computational consumption is also much larger than other approaches.
Appendix C Detailed Experimental Settings
C.1 ComSD Hyper-parameters
Table 2: Hyper-parameter settings of ComSD in unsupervised skill discovery.
Hyper-parameter Setting
Skill vector dimensions $64$
Skill vector space $[0,1]$ continuous
Skill update frequency $50$
State embedding MLP in $f_{\theta_{1}}(\cdot)$ ${\rm dim}(s)\rightarrow 1024\rightarrow 1024\rightarrow 64$
Predictor (MLP) in $f_{\theta_{1}}(\cdot)$ $64\times 2\rightarrow 1024\rightarrow 1024\rightarrow 64$
State encoder activation ${\rm ReLU}$
Skill encoder (MLP) $f_{\theta_{2}}(\cdot)$ $64\rightarrow 1024\rightarrow 1024\rightarrow 64$
Skill encoder activation ${\rm ReLU}$
$\beta$ upper bound $w_{high}$ $2$
$\beta$ lower bound $w_{low}$ 00
$f_{high}$ for walker & quadruped $1$
$f_{low}$ for walker & quadruped 00
Fixed coefficient $\alpha$ for walker $0.25$
Fixed coefficient $\alpha$ for quadruped $31e-$ 3
$f_{high}$ for hopper & cheetah $2/3$
$f_{low}$ for hopper & cheetah $1/3$
Fixed coefficient $\alpha$ for hopper $1.25$
Fixed coefficient $\alpha$ for cheetah $1$
RL backbone algorithm DDPG
Number of pre-training frames 2000000
RL replay buffer size 1000000
Action repeat $1$
Seed (random) frames $4000$
Return discount $0.99$
Number of discounted steps for return $3$
Batch size $1024$
Optimizer Adam
Learning rate $41e-$ 4
Actor network (MLP) ${\rm dim}(s)+64\rightarrow 1024\rightarrow 1024\rightarrow{\rm dim}(a)$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}\rightarrow{\rm Tanh}$
Critic network (MLP) ${\rm dim}(s)+64+{\rm dim}(a)\rightarrow 1024\rightarrow 1024\rightarrow 1$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}$
Agent update frequency $2$
Target critic network EMA $0.01$
Exploration stddev clip $0.3$
Exploration stddev value $0.2$
C.2 Skill Combination Experimental Settings
Table 3: Hyper-parameter settings of skill combination adaptation task.
Hyper-parameter Setting
RL backbone algorithm DDPG
Meta-controller training frames 2000000
RL replay buffer size 1000000
Action repeat $1$
Seed (random) frames $4000$
Return discount $0.99$
Number of discounted steps for return $3$
Batch size $1024$
Optimizer Adam
Learning rate $41e-$ 4
Actor network (MLP) ${\rm dim}(s)\rightarrow 1024\rightarrow 1024\rightarrow 64$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}\rightarrow{\rm Tanh}$
Critic network (MLP) ${\rm dim}(s)+64\rightarrow 1024\rightarrow 1024\rightarrow 1$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}$
Agent update frequency $2$
Target critic network EMA $0.01$
Training stddev clip for meta-controller $0.3$
Training stddev value for meta-controller $0.2$
Eval frequency 10000
Number of Eval episodes 10
Eval stddev value for meta-controller $0.2$
Eval stddev value for pre-trained agent (cheetah) $0.2$
Eval stddev value for pre-trained agent (others) 00
C.3 Skill Finetuning Experimental Settings
Table 4: Hyper-parameter settings of skill finetuning adaptation task.
Hyper-parameter Setting
Fixed target skill for ComSD $(0,0.5,0.5,...,0.5)$
RL backbone algorithm DDPG
Number of finetuning frames 100000
RL replay buffer size 1000000
Action repeat $1$
Seed (random) frames $4000$
Return discount $0.99$
Number of discounted steps for return $3$
Batch size $1024$
Optimizer Adam
Learning rate for walker&quadruped $41e-$ 4
Learning rate for hopper&cheetah $52e-$ 5
Actor network (MLP) ${\rm dim}(s)+64\rightarrow 1024\rightarrow 1024\rightarrow{\rm dim}(a)$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}\rightarrow{\rm Tanh}$
Critic network (MLP) ${\rm dim}(s)+64+{\rm dim}(a)\rightarrow 1024\rightarrow 1024\rightarrow 1$
Actor activation ${\rm layernorm(Tanh)}\rightarrow{\rm ReLU}$
Agent update frequency $2$
Target critic network EMA $0.01$
Training stddev clip $0.3$
Training stddev value $0.2$
Eval frequency 10000
Number of Eval episodes 10
Eval stddev value 00
Appendix D Additional Analysis
D.1 What Skills Do ComSD and Competitive Baselines Discover? & Why does ComSD exhibit better adaptation performance than others?
Figure 6: Visualization for representative behaviors discovered by CIC and APS. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4. Skill AKD ranges from 7 to 10 for CIC and 0 to 2 for APS, which means they can’t discover qualified behaviors at different activity levels.
Figure 7: Visualization for representative behaviors discovered by ComSD. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4. Skill AKD ranges from 0 to 10 for ComSD. The results demonstrate that our ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot.
We provide the skill visualization and corresponding skill AKD of the two most competitive baselines, APS (Liu & Abbeel, 2021a) and CIC (Laskin et al., 2022b), in Figure 6. The visualization and skill AKD of our ComSD are shown in Figure 7. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4.
CIC is able to produce continuous movements of high activity, but it can’t generate behaviors at other activity levels (AKD range of CIC’s skills is 7-10). In addition, CIC suffers from insufficient exploitation, i.e., the generated skills are indistinguishable and homogeneous. CIC’s skills all tend to achieve dynamic flipping, which is consistent with the good initial score on walker flip in skill combination (see Section 4.2). However, the behavioral indistinguishability makes it difficult for meta-controllers to learn ideal combinations for a competitive final score. APS can generate diverse behaviors, but it suffers from lazy exploration and also can’t generate behaviors at different activity levels (AKD range of APS’s skills is 0-2). In general, poor exploration causes poor performance in skill finetuning evaluation. In skill combination, the diversity allows the meta-controller to complete downstream tasks through the combination of unqualified learned skills, which coincides with the upward trend of APS training curves (see Section 4.2). In summary, previous advanced methods can’t provide a good balance between behavioral quality and diversity, thus failing to exhibit competitive results across different downstream adaptation evaluations.
By contrast, our ComSD can produce diverse behaviors at different activity levels (AKD range of ComSD’s skills is 0-10), including flipping, lying down, struggling at different speeds, various postures, and so on. This explains why our ComSD can achieve state-of-the-art adaptation performance across both kinds of downstream tasks while other methods cannot.
D.2 Behavioral Diversity Analysis
In this section, we try to analyze the behavioral diversity of each method. For a skill, we use it to sample 1k states and calculate the Mean State (MS) over all sampled states. MS can partially represent the skill to some extent. For each method, we uniformly sample 41 different walker skills. Over 41 skills, we compute the K-Th-nearest-neighbor MS Distance (KTMSD) and the All-K-nearset-neighbor MS Distance (AKMSD) for skill entropy estimation. KTMSD computes the k-th-nearest-neighbor MS Euclidean distance of each skill and takes the mean, while AKMSD considers all the k-nearest neighbors of each skill and takes the mean distance of MS. (Note that in Section 4.4&Appendix D.1, KTD and AKD are employed for state entropy estimation of one skill, while KTMSD and AKMSD are used for skill entropy estimation of one method in this section.) KTMSD and AKMSD can partially evaluate the skill coverage of one method. In addition, we also calculate the MS Range (MSR) for each method. These metrics are all related to behavioral diversity.
The comparison between all 6 methods on behavioral diversity is shown in Figure 8, demonstrating that ComSD has huge advantages on skill entropy (KTMSD and AKMSD). ComSD is also the most competitive method on MSR. In fact, it’s hard to represent an exploratory skill by only MS (exploratory behaviors have much more visited states than static postures), which puts highly exploratory methods (ComSD and CIC) at a disadvantage on these 3 metrics. In this case, ComSD still obtains state-of-the-art evaluation results, which demonstrates that the behavior set discovered by ComSD is of much higher diversity and coverage than other baselines.
Figure 8: The comparison between all 6 methods on behavioral diversity. ComSD discovers a much more diverse skill set than other baselines.

### v2

@@
-ComSD: Balancing Behavioral Quality and Diversity in Unsupervised Skill Discovery
+Balancing Both Diversity and Exploration in Unsupervised Discovery of Complex Behaviors
 Abstract
-Learning diverse and qualified behaviors for utilization and adaptation without supervision is a key ability of intelligent creatures.
-Ideal unsupervised skill discovery methods are able to produce diverse and qualified skills in the absence of extrinsic reward, while the discovered skill set can efficiently adapt to downstream tasks in various ways.
-Maximizing the Mutual Information (MI) between skills and visited states can achieve ideal skill-conditioned behavior distillation in theory.
-However, it’s difficult for recent advanced methods to well balance behavioral quality (exploration) and diversity (exploitation) in practice, which may be attributed to the unreasonable MI estimation by their rigid intrinsic reward design.
-In this paper, we propose Contrastive multi-objectives Skill Discovery (ComSD) which tries to mitigate the quality-versus-diversity conflict of discovered behaviors through a more reasonable MI estimation and a dynamically weighted intrinsic reward.
-ComSD proposes to employ contrastive learning for a more reasonable estimation of skill-conditioned entropy in MI decomposition.
-In addition, a novel weighting mechanism is proposed to dynamically balance different entropy (in MI decomposition) estimations into a novel multi-objective intrinsic reward, to improve both skill diversity and quality.
-For challenging robot behavior discovery, ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot.
-On numerical evaluations, ComSD exhibits state-of-the-art adaptation performance, significantly outperforming recent advanced skill discovery methods across all skill combination tasks and most skill finetuning tasks.
-Codes will be released at https://github.com/liuxin0824/ComSD.
+This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
+Unsupervised skill discovery seeks to dig out diverse and exploratory skills without extrinsic reward, with the discovered skills efficiently adapting to multiple downstream tasks in various ways.
+However, recent advanced methods struggle to well balance behavioral exploration and diversity, particularly when the agent dynamics are complex and potential skills are hard to discern (e.g., robot behavior discovery).
+In this paper, we propose Contrastive multi-objective Skill Discovery (ComSD) which discovers exploratory and diverse behaviors through a novel intrinsic incentive, named contrastive multi-objective reward.
+It contains a novel diversity reward based on contrastive learning to effectively drive agents to discern existing skills, and a particle-based exploration reward to access and learn new behaviors.
+Moreover, a novel dynamic weighting mechanism between the above two rewards is proposed for diversity-exploration balance, which further improves behavioral quality.
+Extensive experiments and analysis demonstrate that ComSD can generate diverse behaviors at different exploratory levels for complex multi-joint robots, enabling state-of-the-art performance across 32 challenging downstream adaptation tasks, which recent advanced methods cannot.
+Codes will be opened after publication.
+This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
 1 Introduction & Research Background
-Reinforcement Learning (RL) has proven its effectiveness in learning useful task-specific skills in the presence of extrinsic rewards  (Mnih et al., 2015; Levine et al., 2016; Ding et al., 2021; Narvekar et al., 2017; Li et al., 2019).
-The success of unsupervised learning in computer vision (Chen et al., 2020; Caron et al., 2020) and natural language processing (Brown et al., 2020; Devlin et al., 2018) further benefits task-specific RL with complex input (Laskin et al., 2020; Kostrikov et al., 2020; Yarats et al., 2021a; Stooke et al., 2021) by improving representation learning.
-However, the agents trained with lots of efforts are always hard to generalize their knowledge to novel tasks due to the task-specific supervision of extrinsic reward (Stooke et al., 2021).
-In addition to the generalization, intelligent agents should also be able to explore environments and learn different useful behaviors without any extrinsic supervision, like human beings.
-For the reasons above, unsupervised skill discovery is proposed and becomes a novel research hotspot (Gregor et al., 2016; Eysenbach et al., 2018).
-As a branch of unsupervised RL  (Pathak et al., 2017; 2019; Burda et al., 2018), unsupervised skill discovery  (Park et al., 2021; 2023; Strouse et al., 2021) also designs task-agnostic rewards and achieves unsupervised pre-training with these rewards, which guarantees the task-agnostic downstream generalization.
-The main difference is that skill discovery requires extra input as conditions, named skill vectors.
-They aim to discover useful task-agnostic policies that are distinguishable by skill vectors.
-The discovered skills can be implemented for downstream tasks in various ways, including skill finetuning, skill combination, skill imitation, and so on.
-Currently, one of the most popular and effective classes of skill discovery is based on Mutual Information (MI) maximization (Eysenbach et al., 2018; Sharma et al., 2019; Liu & Abbeel, 2021a; Yang et al., 2023).
-Most of them try to design intrinsic rewards to estimate and optimize the MI between skill vectors and visited states.
-With RL maximizing intrinsic reward expectation, correlations between skills and states are distilled, i.e., different useful skills are discovered.
-Recent MI-based methods exhibit considerable results in different fields, including map exploration (Campos et al., 2020), robotic manipulation (Plappert et al., 2018) and video games (Bellemare et al., 2013).
-However, when facing challenging robot locomotion where the space of state and action are continuous, complex, and non-linear, existing advanced methods can’t balance the behavioral quality (exploration) and diversity (exploitation) well.
-Specifically, they can either only learn different static postures of lazy exploration (Eysenbach et al., 2018; Liu & Abbeel, 2021a) or only produce highly dynamic behaviors that are homogeneous and indistinguishable (i.e., insufficient exploitation) (Laskin et al., 2022b), while an ideal robot behavior set should contain diverse behaviors at different activity levels, enabling efficient downstream adaptation in various ways.
-Some recent works have also noticed this issue (Yang et al., 2023), as we have.
-However, they only address it on 2D exploration problems but remain powerless for the difficult multi-joint robot behavior discovery.
-Figure 1: ComSD discovers different useful unsupervised skills through a novel multi-objective intrinsic reward.
-Skill combination and skill finetuning together provide comprehensive skill evaluation, considering both behavioral quality and diversity.
-Our ComSD exhibits state-of-the-art adaptation performance on both kinds of downstream tasks, which recent advanced methods cannot.
-We attribute the above quality-versus-diversity conflict to their unreasonable MI estimation and their rigid intrinsic reward design.
-In this paper, we propose ComSD, which tries to mitigate this conflict by a more reasonable MI estimator and a dynamic weighting algorithm for intrinsic reward design.
-Concretely, ComSD decomposes the MI between state transitions and skill vectors into the negative skill-conditioned entropy and state entropy, designing a novel intrinsic reward for MI estimation and optimization (with RL).
-For conditioned entropy, ComSD proposes to employ the contrastive learning result between skills and corresponding state transitions as a better estimation.
-For state entropy, ComSD follows recent works (Liu & Abbeel, 2021a; Yarats et al., 2021b; Laskin et al., 2022b), choosing a popular particle-based estimation (Liu & Abbeel, 2021b).
-Moreover, a novel Skill-based Multi-objective Weighting (SMW) mechanism is proposed to dynamically balance the above two entropy estimations into a novel multi-objective intrinsic reward, to encourage both exploration and exploitation.
-ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels for challenging multi-joint robots, which recent advanced methods cannot.
-For comprehensive numerical evaluation, two different downstream adaptation tasks are employed: skill combination and skill finetuning.
-Skill finetuning is widely used recently (Laskin et al., 2022b; Yuan et al., 2022; Yang et al., 2023) on URLB (Laskin et al., 2021), where a target skill is chosen from the skill set and further finetuned with extrinsic reward.
-However, it’s one-sided to judge the whole skill set with only one skill, and the process of finetuning also introduces much uncertainty.
-In addition, behavioral diversity is also ignored in skill finetuning.
-To this end, we further employ another skill combination (Eysenbach et al., 2018) to evaluate both behavioral quality and diversity, where the learned skills are frozen and a meta-controller is trained to combine the learned skills for downstream task achievement.
-ComSD outperforms all baselines on skill combination significantly and is competitive with state-of-the-art methods on skill finetuning, as shown in Figure 1.
-Our contributions can be summarized as follows: (i) We propose ComSD, a novel unsupervised skill discovery algorithm that discovers different useful behaviors by maximizing the MI objective between skill vectors and state transitions.
-(ii) ComSD proposes to employ the contrastive learning result between skills and states as a more reasonable conditioned entropy estimation in MI decomposition.
-(iii) A novel weighting algorithm, SMW, is presented to produce a novel multi-objective intrinsic reward that improves skill diversity and quality simultaneously.
-(iv) Comprehensive numerical evaluation and detailed analysis show that ComSD outperforms all recent advanced methods, exhibiting state-of-the-art downstream adaptation ability.
-It can produce a qualified skill set consisting of diverse behaviors for challenging multi-joint robots, which recent advanced methods cannot.
+Under the supervision of extrinsic rewards, reinforcement learning (RL) enables agents to acquire task-specific skills  Mnih et al.
+(2015); Levine et al.
+(2016); Ding et al.
+(2021); Narvekar et al.
+(2017); Li et al.
+(2019).
+In contrast, real intelligent creatures, e.g., humans, are able to learn a number of useful task-agnostic behaviors without any supervision.
+To this end, unsupervised skill discovery is proposed to simulate this process for machines Gregor et al.
+(2016); Eysenbach et al.
+(2018); Park et al.
+(2021, 2023); Strouse et al.
+(2021); Shafiullah and Pinto (2022); Kim et al.
+(2023); Li et al.
+(2023); Zhao et al.
+(2022); Jiang et al.
+(2022).
+This approach aims to generate different useful skills through training a skill-conditioned policy with a designed intrinsic reward.
+The discovered skills should be as exploratory and diverse as possible, which enables efficient adaptation to multiple downstream tasks in various ways, including skill finetuning, skill combination, and so on  Eysenbach et al.
+(2018).
+Thus, both behavioral diversity and exploration are crucial for the discovered skills, and it’s significant to well balance them in unsupervised skill discovery.
+Figure 1: Left: The complex agent dynamics and hard-to-distinguish potential behaviors significantly increase the difficulty of skill discovery.
+First, the difficulty of learning exploratory skills is increased.
+For example, learning to flip for a robot is much harder than reaching a far goal in map exploration due to the different agent dynamics.
+Second, the difficulty of discerning skills (i.e., promoting diversity) is increased.
+For example, reaching different goals can be easily represented by ending states, while dynamic flip and walk cannot be differentiated by ending states.
+In addition, behaviors at different exploratory levels are needed for robots (e.g., postures or different speeds) but are not considered in map.
+Right: ComSD’s pipeline.
+ComSD discovers exploratory and diverse behaviors through a novel contrastive multi-objective reward, showing state-of-the-art adaptation performance on both kinds of downstream tasks, which recent advanced methods cannot.
+Currently, many methods have shown promising results in domains with simple agent dynamics and easily distinguishable potential skills, such as map exploration Campos et al.
+(2020), robotic arm reaching Plappert et al.
+(2018), and video games Bellemare et al.
+(2013).
+However, for more challenging domains where the agent dynamics are complex and the potential behaviors are hard to distinguish, like robot behavior discovery Laskin et al.
+(2021), the skill discovery difficulty is dramatically increased (see Figure 1 left), and recent advanced approaches cannot take both behavioral exploration and diversity into account.
+For example, an ideal robot skill set should contain a variety of behaviors at different activity levels, while recent methods can either only learn different static postures (low exploration)  Eysenbach et al.
+(2018); Liu and Abbeel (2021a) or only produce highly dynamic behaviors that are homogeneous (low diversity)  Laskin et al.
+(2022b).
+Some recent works have also noticed this issue Yang et al.
+(2023), but they only address it on maze exploration problems but remain powerless for more complex robots.
+This is because their intrinsic rewards don’t provide effective incentives to seek and discern far-reaching skills for complex agents, and they also ignore the exploration-diversity conflict (see Section 3.3) when potential behaviors are hard to discern.
+In this paper, we propose Contrastive multi-objective Skill Discovery (ComSD) which seeks exploratory and diverse behaviors through a novel contrastive multi-objective reward.
+It designs strong incentives to respectively encourage diversity and exploration through contrastive learning, with a novel dynamic weight proposed for their balance.
+Concretely, this reward aims to maximize state-skill mutual information (MI), which is decomposed into skill-conditioned entropy (related to diversity) and state entropy (related to exploration).
+For skill-conditioned entropy optimization, we conduct state-skill contrastive learning, utilizing the contrastive results as a novel diversity reward.
+For state entropy maximization, we employ a particle-based estimation as the exploration reward.
+With contrastive learning, the above rewards achieve accurate entropy estimation, thus providing effective encouragement.
+In complex domains with hard-to-discern behaviors, we observe a serious mutually exclusive effect between the above two entropy optimization processes (see Section 3.3).
+Thus, a Skill-based Multi-objective Weighting (SMW) mechanism is proposed to dynamically weight the above rewards into the contrastive multi-objective reward for exploration-diversity balance.
+We evaluate the adaptation ability of discovered behaviors on 16 skill combination tasks and 16 skill finetuning tasks (task description in Section 2.1).
+ComSD enables state-of-the-art adaptation performance compared with five popular advanced methods (Figure 1 right).
+A detailed behavioral quality analysis is also provided, demonstrating that ComSD indeed generates qualified robot behaviors of both exploration and diversity, which recent advanced methods cannot.
+Our contributions can be summarized as follows: (i) We propose ComSD to balance exploration and diversity in complex behavior discovery through a novel contrastive multi-objective reward containing a diversity reward and an exploration reward.
+(ii) We propose to utilize the state-skill contrastive results as a novel diversity reward for an effective diversity incentive.
+(iii) A novel dynamic weighting mechanism (SMW) is presented to weight the above two rewards for diversity-exploration balance, further improving behavioral quality.
+(iv) Extensive experiments and detailed analysis show that ComSD can discover diverse robot behaviors at different exploratory levels, enabling state-of-the-art downstream adaptation ability on 32 challenging tasks.
 2 Preliminaries
 2.1 Problem Definition
-Unsupervised skill discovery algorithms aim at discovering a useful and diverse set of agent behaviors in the absence of extrinsic reward, i.e., training a skill-conditioned policy in a task-agnostic environment.
-Concretely, a reward-free Markov Decision Process (MDP) (Bellman, 1957) is considered and defined as $\mathcal{M}^{free}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the distribution of the next state given the current state and action, $\gamma$ is the discount factor, and $d_{0}$ is the distribution of the initial state.
+Unsupervised skill discovery trains a skill-conditioned policy in a task-agnostic environment.
+Concretely, a reward-free Markov Decision Process (MDP) Bellman (1957) is considered and defined as $\mathcal{M}^{free}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the distribution of the next state given the current state and action, $\gamma$ is the discount factor, and $d_{0}$ is the distribution of the initial state.
 What skill discovery algorithms do is to define an intrinsic reward $r^{intr}$ and augment $\mathcal{M}^{free}$ to a intrinsic-reward MDP $\mathcal{M}^{intr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{intr},\gamma,d_{0})$.
@@
 As an example, we provide the pseudo-code of our ComSD in Appendix A.
-To evaluate the adaption ability of discovered skills (i.e., the skill-conditioned agent $\pi(a|s,z)$), two adaptation evaluations: skill combination and skill finetuning are employed on each task-specific downstream task.
+To evaluate the quality of discovered skills (i.e., the skill-conditioned agent $\pi(a|s,z)$), two adaptation evaluations: skill combination and skill finetuning are employed on each task-specific downstream task.
 A downstream task can be described as an extrinsic-reward MDP $\mathcal{M}^{extr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{extr},\gamma,d_{0})$, where $r^{extr}$ denotes the task-specific extrinsic reward.
@@
 2.2 Mutual Information Objective
-As most skill discovery algorithms do, our ComSD also tries to maximize the Mutual Information (MI) objective between states $\tau(s)$ and skills $z$.
-In general, $\tau(s)$ can be (i) maintaining the original state $\tau(s)=s$, (ii) concatenating the neighboring state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$, or (iii) using the whole trajectory $\tau(s)={\rm concat}(s^{1},...,s^{t})$.
-Following the recent advanced method (Laskin et al., 2022b), we define $\tau$ as state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$ throughout this paper.
+As current state-of-the-art approaches do Laskin et al.
+(2022b); Liu and Abbeel (2021a); Eysenbach et al.
+(2018), our ComSD tries to maximize the Mutual Information (MI) objective between state function $\tau(s)$ and skill vector $z$ for behavior excavation.
+$\tau$ is defined as state transitions ($\tau(s)={\rm concat}(s^{t-1},s^{t})$) throughout this paper.
 The MI objective $I(\tau;z)$ can be decomposed into the form of Shannon entropy in two ways:
-$I(\tau;z)=-H(z|\tau)+H(z)$ (1)
-$I(\tau;z)=-H(\tau|z)+H(\tau)$ (2)
+$I(\tau;z)=-H(z|\tau)+H(z),$ (1)
+$I(\tau;z)=-H(\tau|z)+H(\tau),$ (2)
 where the $I(\cdot;\cdot)$ denotes the MI function and $H(\cdot)$ denotes the Shannon entropy throughout the paper.
-Most classical MI-based algorithms (Gregor et al., 2016; Achiam et al., 2018; Eysenbach et al., 2018; Lee et al., 2019)are based on the first decomposition Eq.
-1.
-A uniform random policy can guarantee the maximum of the skill entropy $H(z)$ while the trainable discriminators estimating the conditioned entropy $H(z|\tau)$ are employed to calculate the intrinsic reward for unsupervised RL.
+Most classical MI-based algorithms Gregor et al.
+(2016); Achiam et al.
+(2018); Eysenbach et al.
+(2018); Lee et al.
+(2019) are based on the first decomposition Eq.
+(1).
+A uniform random policy can guarantee the maximum of the skill entropy $H(z)$ while the trainable discriminators estimating the state-conditioned entropy $H(z|\tau)$ are employed as the intrinsic rewards for unsupervised RL.
 However,  Laskin et al.
-(2021; 2022b) have found it hard for these methods to guarantee robot behavioral activity and exploration without external balancing signals.
-To this end, some recent works (Sharma et al., 2019; Liu & Abbeel, 2021a; Laskin et al., 2022b) focus on the second decomposition Eq.
-2.
-They explicitly optimize the state entropy $H(\tau)$ by exploratory intrinsic rewards, achieving considerable performance on several tasks, including robotic manipulation, video games, and robot locomotion.
+(2021, 2022b) have found it hard for these methods to guarantee robot behavioral exploration without any external balancing signals.
+To this end, some recent works Sharma et al.
+(2019); Liu and Abbeel (2021a); Laskin et al.
+(2022b) focus on the second decomposition Eq.
+(2).
+They explicitly optimize the state entropy $H(\tau)$ by designing exploration rewards, achieving considerable performance in different domains, including robotic manipulation, video games, and challenging robot locomotion.
 Following these, we choose Eq.
-2 as our optimization target.
-3 Unsupervised Skill Discovery by ComSD
-Figure 2: Left: The multi-objective intrinsic reward design of ComSD. Right: In SMW, $\beta$ is linear related to $\rm flag\it(z)$ with different slopes in different region.
-In this section, we detail how ComSD designs a novel multi-objective intrinsic reward based on Eq.
-2.
-In Section 3.1, we propose and describe how to employ contrastive learning to design an exploitation reward for the negative conditioned entropy estimation.
-In Section 3.2, we illustrate the way to estimate the state entropy by a particle-based exploration reward, which follows the recent advanced methods.
-In Section 3.3, we describe why and how to balance the above two sub-intrinsic rewards into a multi-objective intrinsic reward with proposed SMW.
-The overall intrinsic reward design of ComSD is visualized in Figure 2.
-3.1 Conditioned Entropy Estimation via Contrastive Learning
+(2) as our optimization target.
+3 Behavior Discovery by ComSD
+Figure 2: Left: The contrastive multi-objective reward $r^{intr}_{ComSD}$ design.
+Right: In SMW, $\beta$ is linear related to $\rm flag\it(z)$ with different slopes in different region.
+The skill space is divided into different regions with different learning objectives.
+In this section, we show how ComSD seeks complex behaviors via a novel contrastive multi-objective reward $r^{intr}_{ComSD}$ based on Eq.
+(2).
+In Section 3.1, we propose a novel diversity reward $r^{contrast}_{diversity}$ through contrastive learning to estimate the conditioned entropy.
+In Section 3.2, we illustrate the way to estimate the state entropy by a particle-based exploration reward $r_{exploration}$.
+In Section 3.3, we detail why and how to balance the above two rewards into our contrastive multi-objective reward with the proposed SMW.
+The overall design of contrastive multi-objective reward is shown in Figure 2.
+3.1 Skill-conditioned Entropy Estimation via Contrastive Learning
 In decomposition Eq.
-2, increasing the negative conditioned entropy $-H(\tau|z)$ intuitively encourages the agent to access familiar and visited state pairs according to the corresponding skills.
+(2), increasing the negative conditioned entropy $-H(\tau|z)$ intuitively encourages the agent to access familiar and visited state pairs according to the corresponding skills.
 This helps the agent exploit skill-specific information, i.e., increase the diversity of learned skills.
@@
 First, $-H(\tau|z)$ can be decomposed as:
-$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau|z)=\sum_{\tau,z}p(\tau,z)\log\frac{p(\tau,z)}{p(z)}$ (3)
+$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau|z)$ (3)
+$=\sum_{\tau,z}p(\tau,z)\log\frac{p(\tau,z)}{p(z)}$
 $=\sum_{\tau,z}p(\tau,z)(\log p(\tau,z)-\log p(z)),$
@@
 Therefore, the term $\rm log\it p(z)$ becomes a constant, and we denote it as $c$:
-$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)(\rm log\it p(\tau,z)-c)=\sum_{\tau,z}p(\tau,z)\rm log\it p(\tau,z)-\sum_{\tau,z}p(\tau,z)\cdot c$ (4)
-$=\sum_{\tau,z}p(\tau,z)logp(\tau,z)-c\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z).$
-Then we can obtain the variational lower bound:
-$-H(\tau|z)$ $\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z)$ (5)
-$=\sum_{\tau,z}p(\tau,z)\log p(\tau,z)+\sum_{\tau,z}p(\tau,z)\log q(\tau,z)-\sum_{\tau,z}p(\tau,z)\log q(\tau,z)$
-$=\sum_{\tau,z}p(\tau,z)\log q(\tau,z)+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$
-$=\mathbb{E}_{\tau,z}[\log q(\tau,z)]+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))\geq\mathbb{E}_{\tau,z}[\log q(\tau,z)],$
+$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)(\rm log\it p(\tau,z)-c)$ (4)
+$=\sum_{\tau,z}p(\tau,z)\rm log\it p(\tau,z)-\sum_{\tau,z}p(\tau,z)\cdot c$
+$=\sum_{\tau,z}p(\tau,z)logp(\tau,z)-c.$
+We neglect $c$ and then obtain the variational lower bound:
+$-H^{\prime}(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau,z)$ (5)
+$=\sum_{\tau,z}p(\tau,z)\log q(\tau,z)+D_{\rm KL}(p(\tau,z)||q(\tau,z))$
+$=\mathbb{E}_{\tau,z}[\log q(\tau,z)]+D_{\rm KL}(p(\tau,z)||q(\tau,z))$
+$\geq\mathbb{E}_{\tau,z}[\log q(\tau,z)],$
 where $q(\tau,z)$ is employed to estimate the unavailable $p(\tau,z)$.
-To optimize $-H(\tau|z)$, we need to minimize the KL divergence $D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$ and maximize the expectation $\mathbb{E}_{\tau,z}[\log q(\tau,z)]$.
+To optimize $-H(\tau|z)$, we need to minimize the KL divergence $D_{\rm KL}(p(\tau,z)||q(\tau,z))$ and maximize the expectation $\mathbb{E}_{\tau,z}[\log q(\tau,z)]$.
 For KL divergence minimization, the distribution $q$ should be as positively correlated with the probability function $p$ as possible.
@@
 where $f_{\theta_{1}}(\cdot)$ is the state encoder, $f_{\theta_{2}}(\cdot)$ is the skill encoder, T is the temperature, and $\exp(\cdot)$ makes $q$ non-negative like $p$.
-The above two encoders are updated by gradient descent after computing the NCE loss (Gutmann & Hyvärinen, 2010) in contrastive learning (Chen et al., 2020):
+The above two encoders are updated by gradient descent after computing the NCE loss Gutmann and Hyvärinen (2010) in contrastive learning Chen et al.
+(2020):
 $\mathcal{L}_{NCE}=\log q(\tau_{i},z_{i})-\log\frac{1}{N}\sum_{j=1}^{N}q(\tau_{j},z_{i}),$ (7)
@@
 Maximizing $\mathcal{L}_{NCE}$ increases the value of $q(\tau_{i},z_{i})$ and decreases the value of $q(\tau_{j},z_{i})$, which actually shrinks the difference between $p$ and $q$.
-For expectation maximization, we regard the optimization objective as a part of our unsupervised intrinsic reward, employing RL to optimize it:
-$r^{intr}_{exploitation}\propto q(\tau_{i},z_{i}).$ (8)
+For expectation maximization, we regard the content (the contrastive result) as our novel contrastive diversity reward, employing RL to optimize it:
+$r^{contrast}_{diversity}\propto q(\tau_{i},z_{i}).$ (8)
 3.2 Particle-based State Entropy Estimation
 In decomposition Eq.
-2, increasing the state entropy $H(\tau)$ encourages the agent to explore more widely and visit more state transitions.
-Following  Liu & Abbeel (2021a); Yarats et al.
+(2), increasing the state entropy $H(\tau)$ encourages the agent to explore more widely and visit more state transitions.
+Following  Liu and Abbeel (2021a); Yarats et al.
 (2021b); Laskin et al.
-(2022b), we employ a particle-based entropy estimation proposed by  Liu & Abbeel (2021b).
-Concretely, $H(\tau)$ can be estimated by the Euclidean distance between each particle ($\tau$ in our paper) and its all $k$-nearest neighbor (Singh et al., 2003) in the latent space:
+(2022b), we employ a particle-based entropy estimation proposed by  Liu and Abbeel (2021b).
+Concretely, $H(\tau)$ can be estimated by the Euclidean distance between each particle ($\tau$ in our paper) and its all $k$-nearest neighbor Singh et al.
+(2003) in the latent space:
 $H(\tau)\approx H_{particle}(\tau)\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{buffer}^{k-nn}(h_{i})}\log||h_{i}-h_{i}^{nn}||,$ (9)
@@
 The exploration reward is defined as the following:
-$r^{intr}_{exploration}\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{batch}^{k-nn}(h_{i})}\log||h_{i}-h_{i}^{nn}||,\ \ where\ \ h_{i}=f_{\theta_{1}}(\tau_{i}).$ (10)
-3.3 Skill-based Multi-objectives weighting
-With the negative conditioned entropy $-H(\tau|z)$ estimated by $r^{intr}_{exploitation}$ and the state entropy estimated by $r^{intr}_{exploration}$, a naive way for intrinsic reward design is to employ a fixed coefficient $\alpha$ to scale the two terms ($r^{intr}_{naive}=r^{intr}_{exploration}+\alpha r^{intr}_{exploitation}$).
-However, we find it impossible to choose a proper fixed coefficient to balance them well especially in challenging multi-joint robot behavior discovery, which is because of a severe quality-versus-diversity conflict between the above two sub-intrinsic rewards.
-Specifically, when employing $r^{intr}_{exploration}$ alone, the agents can only learn dynamic behaviors of high activity but not easing movements or static postures, and the learned skills are highly homogeneous and indistinguishable by skill vectors (i.e., unsatisfactory exploitation).
-The naive intervention of $r^{intr}_{exploitation}$ can significantly improve the behavioral diversity, but it will simultaneously cause the lazy exploration, making the robots unable to discover active behaviors.
-We propose Skill-based Multi-objectives Weighting (SMW), a simple but effective dynamic weighting mechanism for the issue above.
-SMW aims to take advantage of the two sub-rewards by setting different optimization objectives for different skill vectors.
-Concretely, it designs another changeable coefficient $\beta$ which is dynamically adjusted according to skill vectors.
+$r_{exploration}\propto\frac{1}{k}\sum_{h_{i}^{nn}\in N_{batch}^{k-nn}(h_{i})}\log||h_{i}$ $-h_{i}^{nn}||,$ (10)
+$where\ \ h_{i}$ $=f_{\theta_{1}}(\tau_{i}).$
+Figure 3: The training curve of all 6 methods on 16 downstream skill combination tasks.
+ComSD outperforms all baselines significantly across all 16 tasks, demonstrating that ComSD discovers much more exploratory and diverse behaviors than other methods for challenging multi-joint robots.
+3.3 Skill-based Multi-objective Weighting
+With the negative conditioned entropy $-H(\tau|z)$ estimated by $r^{contrast}_{diversity}$ and the state entropy estimated by $r_{exploration}$, a naive way for intrinsic reward design is to employ a fixed coefficient $\alpha$ to scale the two terms ($r^{intr}_{naive}=r_{exploration}+\alpha\cdot r^{contrast}_{diversity}$).
+However, in practice, we find it impossible to use a fixed coefficient to balance them well when potential behaviors are complex and hard to distinguish by only ending states.
+Concretely, when employing $r_{exploration}$ alone, the robots can only learn dynamic behaviors of high activity but not other exploratory levels, and the learned behaviors are highly homogeneous.
+Increasing the fixed weight $\alpha$ of $r^{contrast}_{diversity}$ (or other $-H(\tau|z)$ estimation in other methods) can surely improve the behavioral diversity, but it will simultaneously reduce the exploration (see experiments in Figure 6), making robots prefer static postures (distinguishable by ending states) and ignore active behaviors (indistinguishable by ending states).
+This is because distilling ending states into skill vectors is an easy-to-reach local optimal solution for state-skill MI maximization.
+We propose Skill-based Multi-objective Weighting (SMW), an easy-but-effective weighting mechanism for the issue above.
+The main idea of SMW is to set different learning objectives for different skills instead of using fixed learning objectives across all skills.
+Concretely, it designs another changeable coefficient $\beta(z)$ which is dynamically adjusted according to skill vectors.
 $\beta(z)$ is defined as:
@@
 […]
-The $\rm flag\it(\cdot)$ can be any function to map the high-dimensional skill vectors into 1-dimensional flag variables.
-We find that simply employing the first dimension of skill vectors (i.e., $\rm flag\it(z)=I_{1}^{T}\cdot z,\ {\rm where}\ I_{1}=(1,0,...,0)$) can perform well.
-The fixed hyper-parameters $w_{high},w_{low},f_{high},f_{low}$ and the clamp function $\rm clamp\it(\cdot)$ make the dynamic coefficient $\beta$ linearly related to flag variables $\rm flag\it(z)$ with different slopes in different regions.
-We clearly visualize the relation between $\beta$ and $\rm flag\it(z)$ in Figure 2.
-With SMW, the multi-objective intrinsic reward of ComSD is defined as:
-$r^{intr}_{ComSD}=r^{intr}_{exploration}+\alpha\cdot\beta(z)\cdot r^{intr}_{exploitation}.$ (12)
-Over the intrinsic-reward MDP $\mathcal{M}^{intr}$, the skill-conditioned policy $\pi(a|s,z)$ can be trained by RL with $z\sim p(z)$. Algorithm 1 in Appendix A provides the full pseudo-code of ComSD.
-4 Skill Evaluation and Analysis
+[…]
+$r^{intr}_{ComSD}=r_{exploration}+\alpha\cdot\beta(z)\cdot r^{contrast}_{diversity}.$ (12)
+Over the intrinsic-reward MDP $\mathcal{M}^{intr}$ augmented by $r^{intr}_{ComSD}$, ComSD trains the skill-conditioned policy $\pi(a|s,z)$ via RL with $z$ sampled by uniform distribution.
+Appendix A provides the full pseudo-code of ComSD.
+4 Experiments: Skill Adaptation and Analysis
 4.1 Experimental Setup
 Environments.
-On the importance of environments: OpenAI Gym (Brockman et al., 2016) and DMControl (Tassa et al., 2018) are two most popular continuous robot locomotion benchmarks.
-In Gym, the episode is ended when agents lose their balance, while the episode length in DMControl is fixed.
+On the importance of environments: OpenAI Gym Brockman et al.
+(2016) and DMC Tassa et al.
+(2018) are two most popular continuous robot locomotion benchmarks.
+In Gym, the episode is ended when agents lose their balance, while the episode length in DMC is fixed.
 Laskin et al.
-(2022b) found this difference makes DMControl much harder for reward-free exploration since agents have to learn balance by themselves without any external signals.
-Following recent advanced methods (Laskin et al., 2022b; Yarats et al., 2021b; Liu & Abbeel, 2021a; b; Liu et al., 2023), we employ 16 downstream tasks of 4 domains from URLB (Laskin et al., 2021) and DMControl (Tassa et al., 2018) for skill evaluation.
-The domains Walker, Quadruped, Cheetah and Hopper are recognized as the most representative and challenging multi-joint robot environments, each of which contains 4 totally diverse and challenging downstream locomotion tasks.
-Evaluations.
-In each domain, all the methods, including ComSD pre-train their agents for 2M environment steps with their respective intrinsic rewards.
-After unsupervised pre-training, the behavioral diversity and quality will be fully and reasonably evaluated on two adaptation tasks: skill combination (Eysenbach et al., 2018) and skill combination in URLB (Laskin et al., 2021) respectively, across all 16 downstream tasks with extrinsic reward.
-It means a total of 32 numerical results are employed for one method’s evaluation.
-DDPG (Lillicrap et al., 2015) is chosen as the backbone RL algorithm for all methods throughout the paper.
-The detailed settings of the skill combination and skill finetuning are provided along with the experimental results in Section 4.2.
-Baselines.
-We compare our ComSD with five recent advanced skill discovery algorithms that are popular for robot behavior discovery.
-They are BeCL (Yang et al., 2023), CIC (Laskin et al., 2022b), APS (Liu & Abbeel, 2021a), SMM (Lee et al., 2019), and DIAYN (Eysenbach et al., 2018).
-All the methods try to optimize their MI target by designing intrinsic rewards for unsupervised RL.
-In addition to BeCL, they all try to maximize $I(\tau,z)$, where DIAYN and SMM choose the decomposition Eq.
-1 while APS, CIC, and our ComSD choose decomposition Eq.
-2.
-BeCL proposes a novel MI objective $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill.
-We provide a detailed description of these baselines in Appendix B.
-The close baselines to our ComSD are CIC and APS.
-Apart from SMW, a novel weighting mechanism, ComSD still differs significantly from these methods.
-For CIC, ComSD follows it in state entropy estimation but first proposes to employ contrastive results for explicit state entropy maximization.
-APS and ComSD both explicitly optimize the state entropy and conditioned entropy, while ComSD utilizes contrastive learning for a more reasonable estimation.
-In addition, BeCL also employs contrastive learning like ComSD, but their contrastive components are different.
-4.2 Main Results
-Figure 3: The training curve of all 6 methods on 16 downstream skill combination tasks.
-ComSD outperforms all baselines significantly across all 16 tasks, demonstrating that ComSD discovers much more diverse and qualified behaviors than other methods for challenging multi-joint robots.
-Skill Combination.
-In skill combination, we train another meta-controller $\pi^{\prime}(z|s)$ which selects the discovered skills automatically to achieve downstream tasks.
-The discovered skill-conditioned agent $\pi(a|s,z)$ is frozen, and $\pi^{\prime}(z|s)$ is trained by RL with extrinsic reward (see Section 2.1 and Appendix A.2 for a detailed problem definition).
-This adaptation task can effectively evaluate both the diversity and quality of the discovered skill set.
-For each method, 2M environment steps RL are allowed to train their meta-controllers.
-All the hyper-parameters of meta-controller and RL settings are shared across all the methods for a fair comparison, which we detail in Appendix C.
-The results of ComSD are obtained over 6 different random seeds.
-The training curves are shown in Figure 3, which demonstrates that ComSD significantly outperforms all five baselines across all 16 downstream tasks.
-In skill combination, the better adaptation performance means that the behaviors discovered by ComSD are of higher diversity and quality than other methods.
-CIC and APS are the two most competitive baselines.
-Both of them and our ComSD employ the decomposition Eq.
-2, which indicates the importance of explicitly maximizing state entropy for discovering robot behaviors.
-An interesting phenomenon is that CIC always gets a higher initial score than APS and ComSD but has no upward trend, which coincides with the fact that CIC can only find homogeneous behaviors at a high activity level.
-By contrast, APS and ComSD explicitly maximize the conditioned entropy through intrinsic rewards, greatly improving the diversity of learned skill sets and achieving higher final scores than CIC.
-Compared with APS, ComSD performs better due to the following two reasons: First, ComSD employs contrastive learning for a better estimation of both state entropy and conditioned entropy.
-Second, the proposed SMW enables ComSD to learn qualified behaviors at different activity levels, thus gaining better adaptation ability.
-Skill Finetuning.
-In skill finetuning, a skill vector $z_{i}$ is chosen, and the corresponding skill $\pi(\cdot|s,z_{i})$ is finetuned with the extrinsic reward (see Section 2.1 and Appendix A.3 for a detailed problem definition).
-100k environment steps with extrinsic reward are allowed for each method, where a proper skill vector is selected in the first 4k steps and the chosen skill is finetuned in another 96k steps.
-For skill selection, previous works employ random sampling (Yang et al., 2023), fixed choice (Laskin et al., 2022b; a), or reward-based choice (Liu & Abbeel, 2021a).
-We follow the official implementation of  Laskin et al.
-(2022b), employing a fixed mean skill with the first dimension set to 00.
-All the hyper-parameters of neural architecture and RL are shared across different methods for a fair comparison.
-We show the detailed experimental settings in Appendix C.
-The results of ComSD are obtained over 10 different random seeds.
-Table 1: The numerical results of all 6 methods on 16 skill finetuning downstream tasks.
-ComSD exhibits competitive adaptation performance compared with state-of-the-art methods (CIC and BeCL).
-In skill finetuning, only one skill is used for evaluation.
-The behavioral diversity is also ignored.
-It is one-sided to employ skill finetuning alone for evaluation, like recent works.
+(2022b) found this difference makes DMC much harder for reward-free exploration since agents have to learn balance by themselves without any external signals.
+Following recent advanced methods Laskin et al.
+(2022b); Yarats et al.
+(2021b); Liu et al.
+(2023); Laskin et al.
+(2020); Stooke et al.
+(2021); Yarats et al.
+(2021a), we employ 16 downstream tasks of 4 domains from URLB Laskin et al.
+(2021) and DMC Tassa et al.
+(2018) for skill adaptation evaluation.
+The domains Walker, Quadruped, Cheetah, and Hopper are recognized as the most representative and challenging multi-joint robot environments, each of which contains four totally diverse and challenging downstream tasks.
 Domain Task DIAYN SMM APS CIC BeCL ComSD(ours)
@@
 flip backward 477±108 679±7 518±103 678±93 643±102 685±55
+Table 1: The numerical results of all 6 methods on 16 skill finetuning downstream tasks. ComSD exhibits competitive adaptation performance compared with state-of-the-art methods (CIC and BeCL).
+Evaluations.
+In each domain, all the methods pre-train their agents for 2M environment steps with their respective intrinsic rewards.
+After unsupervised pre-training, the behavioral quality will be fully evaluated on two adaptation tasks: skill combination Eysenbach et al.
+(2018) and skill combination in URLB Laskin et al.
+(2021) respectively, across all 16 downstream tasks with extrinsic reward.
+It means a total of 32 numerical results are employed for one method’s evaluation.
+DDPG Lillicrap et al.
+(2015) is chosen as the backbone RL algorithm for all methods throughout the paper.
+The detailed settings of the skill combination and skill finetuning are provided along with the experimental results in Section 4.2.
+The direct quality analysis is also provided in Section 4.3.
+Baselines.
+We compare our ComSD with five recent advanced algorithms for robot behavior discovery.
+They are BeCL Yang et al.
+(2023), CIC Laskin et al.
+(2022b), APS Liu and Abbeel (2021a), SMM Lee et al.
+(2019), and DIAYN Eysenbach et al.
+(2018).
+We provide detailed description of these baselines in Appendix B.
+The close baselines to our ComSD are CIC and APS, where they all design intrinsic rewards based on Eq.
+(2).
+Apart from the most notable difference, a novel weighting mechanism (SMW), ComSD still significantly differs from these methods.
+CIC doesn’t design a diversity reward, while ComSD proposes a novel contrastive diversity reward to explicitly encourage diversity.
+APS also employs both diversity reward and exploration reward, but it doesn’t conduct contrastive learning which is the key to provide effective incentives for complex agents.
+In addition, BeCL also employs contrastive learning like us, but the contrastive components are totally different.
+Figure 4: Adaptation ablation experiments on (left) two skill combination tasks and (right) two skill finetuning tasks.
+Contrastive diversity reward and SMW are both necessary for ComSD to achieve the most competitive results on both kinds of adaptation tasks.
+Figure 5: Left three: Behavioral exploration analysis of all methods.
+AKD can represent the exploratory level of one skill.
+ComSD leads on all three metrics, which means it generates behaviors at different exploratory levels, including dynamic movements and static postures, which other methods cannot.
+Right one: Behavioral diversity analysis of all methods.
+ComSD discovers the most diverse robot behaviors.
+4.2 Adaptation Results
+Skill Combination.
+In skill combination, we train another meta-controller $\pi^{\prime}(z|s)$ which selects the discovered skills automatically to achieve downstream tasks.
+The discovered skill-conditioned agent $\pi(a|s,z)$ is frozen, and $\pi^{\prime}(z|s)$ is trained by RL with extrinsic reward (see Section 2.1 and Appendix A for a detailed problem definition).
+This adaptation task can effectively evaluate behavioral quality (both exploration and diversity).
+For each method, 2M environment steps RL are allowed to train their meta-controllers.
+All the hyper-parameters are shared across all the methods for a fair comparison, which we detail in Appendix C.
+The results of ComSD are obtained over 6 different random seeds.
+The training curves are shown in Figure 3, which demonstrates that ComSD significantly outperforms all five baselines across all 16 downstream tasks.
+In skill combination, the better adaptation performance means that the behaviors discovered by ComSD are of higher quality (both exploration and diversity) than other methods.
+CIC and APS are the two most competitive baselines.
+Both of them and our ComSD employ the decomposition Eq.
+(2), which indicates the importance of explicitly maximize exploration.
+An interesting phenomenon is that CIC always gets a higher initial score than APS and ComSD but has no upward trend, which coincides with the fact that CIC can only find homogeneous behaviors at a high exploratory level.
+By contrast, APS and ComSD explicitly maximize the conditioned entropy through diversity reward, greatly improving the diversity of learned skill sets and achieving higher final scores than CIC.
+Compared with APS, ComSD performs better due to the following two reasons: First, the contrastive learning employed by ComSD provides stronger incentives for skill exploration and exploitation in complex domains.
+Second, the proposed SMW enables ComSD to learn diverse behaviors at different activity levels, thus gaining better multi-task adaptation.
+Skill Finetuning In skill finetuning, a skill vector $z_{i}$ is chosen, and the corresponding skill $\pi(\cdot|s,z_{i})$ is finetuned with the extrinsic reward (see Section 2.1 and Appendix A for a detailed problem definition).
+100k environment steps with extrinsic reward are allowed for each method, where one skill is selected in the first 4k steps and then finetuned in another 96k steps.
+For skill selection, previous works employ random sampling Yang et al.
+(2023), fixed choice Laskin et al.
+(2022b, a), or reward-based choice Liu and Abbeel (2021a).
+We follow the official implementation of  Laskin et al.
+(2022b), employing a fixed mean skill with the first dimension set to 00.
+All the hyper-parameters are shared across different methods for a fair comparison.
+We show the detailed experimental settings in Appendix C.
+The results of ComSD are obtained over 10 different random seeds.
 The numerical results on 16 downstream tasks are shown in Table 1.
@@
 This explains why CIC and BeCL use random or fixed choices.
-Furthermore, higher behavioral diversity causes higher choice difficulty in ComSD.
-ComSD discovers the most diverse behaviors (state-of-the-art adaptation on skill combination) while simultaneously achieving competitive performance on skill finetuning with a fixed skill, which further shows its superior behavioral quality.
-Figure 4: Numerical ablation experiments on (left) two skill combination tasks and (right) two skill finetuning tasks.
-Contrastive estimation of negative conditioned entropy and SMW are both necessary for ComSD to achieve the most competitive results on both kinds of adaptation tasks.
-Figure 5: Skill activity analysis of our ComSD, ComSD w/o SMW, and all baselines.
-We estimate the skill activity by visited state entropy (KTD and AKD) and report the activity variance and range of uniformly selected walker skills for each method.
-ComSD exhibits much higher entropy variance and range, demonstrating that ComSD indeed discovers skills at different activity levels.
-4.3 Numerical Ablation Study
-In this section, we conduct numerical ablation experiments to show the effectiveness of both contrastive conditioned entropy estimation and SMW.
-The results are shown in Figure 4, where CIC is actually ComSD w/o explicit conditioned entropy estimation&SMW, and APS is actually ComSD w/o contrastive conditioned entropy estimation&SMW.
-Compared with implicit maximization (CIC), explicitly maximizing the skill-conditioned entropy (APS and ComSD w/o SMW) improves the training curve slope on skill combination, which means better exploitation, i.e., behavioral diversity.
-Moreover, the contrastive result is a better conditioned entropy estimation, improving APS on both kinds of adaptation tasks.
-However, simply maximizing the conditioned entropy explicitly leads to a trivial solution in practice, where the agent gives up continuous movement and exploration.
-They can only learn static postures but not dynamic behaviors, exhibiting performance drops on skill finetuning compared with CIC.
-This phenomenon is also observed by Laskin et al.
-(2022b).
-SMW effectively alleviates this exploration hurt by setting multiple optimization objectives.
-It improves behavioral exploitation while maintaining the tendency toward exploration, making ComSD exhibit state-of-the-art performance on both kinds of adaptation tasks.
-In summary, SMW and contrastive conditioned entropy estimation are both necessary for ComSD.
-4.4 Quality Analysis of Learned Skills
-Numerical results on 32 downstream tasks have demonstrated the effectiveness and superiority of ComSD.
-In this section, we directly analyze the learned skills to show that our method indeed discovers behaviors at different activity levels, which previous advanced methods cannot.
-For each method, we uniformly select eleven walker skills and use the entropy of visited states to represent the skill activity.
-K-Th-nearest-neighbor Distance (KTD) and All-K-nearest-neighbor Distance (AKD) (Liu & Abbeel, 2021b), are employed as state entropy estimations.
-KTD computes the k-th-nearest-neighbor Euclidean distance of each state and takes the mean, while AKD considers all the k-nearest neighbors of each state and takes the mean distance.
-For each method, we report the entropy variance and entropy range of the selected skills in Figure 5.
-ComSD discovers a skill set with the biggest activity range and variance, demonstrating it indeed produces skills at diverse activity levels, where mainly attributes to multiple optimization targets brought by multi-objective intrinsic reward.
-We strongly refer readers to Appendix D for additional visualization and analysis.
+Furthermore, higher behavioral diversity causes higher choice difficulties.
+ComSD discovers the most diverse behaviors (see skill combination results and analysis in section 4.3) while simultaneously achieving competitive performance on skill finetuning, which further shows its superior skill quality.
+Adaptation Ablation.
+We conduct ablation experiments on both kinds of adaptation tasks to show the effectiveness of (i) contrastive diversity reward $r^{contrast}_{diversity}$ and (ii) SMW in Figure 4, where CIC is actually ComSD w/o $r^{contrast}_{diversity}$&SMW, and APS is actually ComSD w/o contrastive learning&SMW.
+Compared with no diversity reward (CIC), designing diversity reward (APS and ComSD w/o SMW) enables better behavioral diversity, corresponding to larger curve slope on skill combination.
+Moreover, contrastive learning enables ComSD w/o SMW to outperform APS on both kinds of adaptation tasks.
+In practice, a fixed diversity reward (APS and ComSD w/o SMW) reduces the behavioral exploration (see Section 4.3).
+This causes severe performance drops on skill finetuning, which is effectively alleviated by SMW.
+In summary, both $r^{contrast}_{diversity}$ and SMW are needed for ideal adaptation.
+4.3 Behavioral Quality Analysis
+Behavioral Exploration.
+In this section, we analyze the behavioral exploration (activity) of all methods.
+We use the visited-state coverage (entropy) of one skill to represent its exploration.
+The visited state entropy is estimated by the All-K-nearest-neighbor Distance (AKD) Liu and Abbeel (2021b) that computes the Euclidean distance between each state and its k nearest neighbors, and then takes the mean.
+AKD can denote the exploratory level of one skill.
+For each method, we use grid sweep  Laskin et al.
+(2022b) to select eleven walker skills and report the variance, range, and maximum of AKD in Figure 5.
+ComSD and CIC lead on AKD-max, which means they both discover the most exploratory skill while other methods cannot.
+In addition, ComSD exhibits huge superiority in skill activity range and variance, indeed discovering behaviors at different exploratory levels, which all other methods cannot.
+Behavioral Diversity.
+In this section, we try to analyze the behavioral diversity of all methods.
+For one skill, we represent it by the Mean State (MS) over 1k visited states.
+For each method, we use grid sweep  Laskin et al.
+(2022b) to select several walker behaviors, and then calculate the MS-coverage by k-nearest-neighbor Liu and Abbeel (2021b) Euclidean distance.
+MS-coverage can partially represent behavioral diversity to some extent.
+The results in Figure 5 demonstrate ComSD has huge advantages on MS-coverage, discovering the most diverse skill set.
+SMW Moltivation & Behavioral Quality Ablation.
+We ablate SMW, setting different fixed coefficients $\alpha$ of $r^{contrast}_{diversity}$ to show the effect on behavioral quality.
+As shown in Figure 6 left, there exists a serious conflict between exploration and diversity.
+Concretely, $r^{contrast}_{diversity}$ simultaneously improves behavioral diversity and hurts exploration, and so do other diversity incentives in other methods (e.g., APS).
+Any fixed weight $\alpha$ cannot achieve diversity-exploration balance.
+This motivates our SMW which links the weight to the changing skill vectors, achieving dynamic weighting, as mentioned in Section 3.3.
+In addition, SMW further enables behavior discovery at different exploratory levels (Figure 6 right), which a fixed coefficient also cannot support.
+Skill Visualization.
+We visualize the behaviors of ComSD, CIC, and APS in Appendix D, proving that (i) CIC can only learn dynamic but homogeneous flipping behaviors; (ii) APS acquires different static postures of lazy exploration; and (iii) ComSD discovers diverse behaviors at different exploratory levels, including flipping, standing, falling down, and different yoga postures.
+ComSD’s qualified behaviors enable state-of-the-art performance on different adaptation tasks.
+Figure 6: SMW motivation & Behavioral quality ablation.
+Left: The exploration-diversity conflict in complex domains.
+Increasing the fixed weight $\alpha$ of diversity reward can surely improve the behavioral diversity, but it will simultaneously reduce the skill exploration.
+The dynamic SMW can effectively achieve diversity-exploration balance.
+Right: SMW enables behaviors at different exploratory levels, further improving behavioral quality.
 5 Conclusion
-In this paper, we propose a novel unsupervised skill discovery method named ComSD.
-It improves both behavioral quality and diversity through contrastive learning from a multi-objective perspective.
-For numerical evaluation, we conduct comprehensive comparisons on different kinds of downstream adaptation tasks, considering behavioral diversity ignored by recent works.
-The superiority of ComSD over all other baselines is well verified by our employed evaluation.
-We hope (i) our ComSD can inspire more attention to intrinsic reward weighting algorithms, and (ii) our experimental settings can make a difference for unsupervised skill evaluation in future works.
+In this paper, we propose ComSD to balance behavioral exploration and diversity through contrastive learning from a multi-objective perspective.
+We conduct extensive experiments and detailed analysis, verifying the superiority of ComSD on both skill quality and adaptation ability over all other baselines.
+We hope (i) our ComSD can inspire more attention to the exploration-diversity trade-off in complex behavior discovery, and (ii) our comprehensive experiments can inspire unsupervised skill evaluation for future works.
 Appendix
@@
 Output: The performance of $\pi(a|s,z_{i})$ serves as the skill finetuning evaluation result.
-Appendix B Baseline Details
+Appendix B Extended Backgrounds
+B.1 Related Works
+Reinforcement Learning (RL) has proven its effectiveness in learning useful task-specific skills in the presence of extrinsic rewards  Mnih et al.
+[2015]; Levine et al.
+[2016]; Ding et al.
+[2021]; Narvekar et al.
+[2017]; Li et al.
+[2019].
+The success of unsupervised learning in computer vision Chen et al.
+[2020]; Caron et al.
+[2020] and natural language processing Brown et al.
+[2020]; Devlin et al.
+[2018] further benefits task-specific RL with complex input Laskin et al.
+[2020]; Kostrikov et al.
+[2020]; Yarats et al.
+[2021a]; Stooke et al.
+[2021] by improving representation learning.
+However, the agents trained with lots of efforts are always hard to generalize their knowledge to novel tasks due to the task-specific supervision of extrinsic reward Stooke et al.
+[2021].
+Unsupervised RL  Laskin et al.
+[2021] is proposed to improve this generalization, where the task-agnostic reward is designed for unsupervised pre-training.
+The pre-trained feature encoders  Liu and Abbeel [2021b]; Yarats et al.
+[2021b]; Liu et al.
+[2023] and exploration policies  Pathak et al.
+[2017, 2019]; Burda et al.
+[2018] then can be implemented for efficient RL on different downstream tasks.
+In addition to the generalization, intelligent agents should also be able to explore environments and learn different useful behaviors without any extrinsic supervision, like human beings.
+For the reasons above, unsupervised skill discovery  Gregor et al.
+[2016]; Eysenbach et al.
+[2018]; Park et al.
+[2021, 2023]; Strouse et al.
+[2021]; Shafiullah and Pinto [2022]; Kim et al.
+[2023]; Li et al.
+[2023]; Zhao et al.
+[2022]; Jiang et al.
+[2022]; Liu and Abbeel [2021a]; Lee et al.
+[2019]; Laskin et al.
+[2022b] is proposed and becomes a novel research hotspot.
+As a branch of unsupervised RL, unsupervised skill discovery also designs task-agnostic rewards and achieves unsupervised pre-training with these rewards, which guarantees task-agnostic downstream generalization.
+The main difference is that skill discovery requires extra input as conditions, named skill vectors.
+They aim to discover useful task-agnostic policies that are distinguishable by skill vectors.
+These approaches exhibit promising results across different fields  Brockman et al.
+[2016]; Bellemare et al.
+[2013]; Plappert et al.
+[2018]; Campos et al.
+[2020].
+However, it’s hard for recent advanced methods to take both behavioral exploration and diversity into account, especially when faced with complex domains Laskin et al.
+[2021]; Tassa et al.
+[2018].
+This is because their intrinsic rewards don’t provide effective incentives to seek and discern far-reaching skills for complex agents, and they also ignore the exploration-diversity conflict (see Section 3.3) when potential behaviors are hard to discern.
+By contrast, our ComSD designs strong incentives to respectively encourage diversity and exploration through contrastive learning, with a novel dynamic weight proposed for their balance.
+B.2 Baseline Details
 DIAYN
-(Eysenbach et al., 2018) is one of the most classical and original unsupervised skill discovery algorithms, trying to maximize the MI between skills and states.
+Eysenbach et al.
+[2018] is one of the most classical and original unsupervised skill discovery algorithms, trying to maximize the MI between skills and states.
 It employs the first MI decomposition, Eq.
-1.
+(1).
 It uses a discrete uniform prior distribution to guarantee the maximization of skill entropy $H(z)$.
@@
 SMM
-(Lee et al., 2019) aims to learn a policy for which the state marginal distribution matches a given target state distribution.
+Lee et al.
+[2019] aims to learn a policy for which the state marginal distribution matches a given target state distribution.
 It optimizes the objective by reducing it to a two-player, zero-sum game between a state density model and a parametric policy.
 Like DIAYN, it is also based on the first decomposition (Eq.
-1) of MI and employs discriminator training.
-The difference is that SMM explicitly maximizes the state entropy with intrinsic reward, which inspires lots of recent advanced works and our ComSD.
+(1)) of MI and employs discriminator.
+The difference is that SMM explicitly maximizes state entropy with intrinsic reward, which inspires lots of recent works including ours.
 APS
-(Liu & Abbeel, 2021a) first employs the second MI decomposition Eq.
-2 for a better MI estimation.
-For state entropy estimation, it employs a popular particle-based entropy estimation proposed by APT (Liu & Abbeel, 2021b), which is proven effective and supports many advanced works (Laskin et al., 2022b; Yarats et al., 2021b) and our ComSD.
-For skill-conditioned entropy, it chooses the successor feature (Hansen et al., 2019), introducing it into the final intrinsic reward for an explicit maximization.
-The weight between different entropy estimations is fixed in APS.
-APS can’t guarantee the behavioral quality (exploration) well.
-Different from APS, our ComSD employs contrastive learning for better conditioned entropy estimation and designs a novel dynamic weighting algorithm (SMW) to overcome the exploration drop brought by explicit conditioned entropy maximization.
+Liu and Abbeel [2021a] employs the second MI decomposition Eq.
+(2) for a better MI estimation.
+For state entropy estimation, it employs a popular particle-based entropy estimation proposed by APT Liu and Abbeel [2021b] for exploration, which is proven effective and supports many advanced works Laskin et al.
+[2022b]; Yarats et al.
+[2021b] and our ComSD.
+For skill-conditioned entropy, it chooses the successor feature Hansen et al.
+[2019] as a diversity reward with a fixed weight.
+However, we find that APS’s reward can’t provide enough incentives for complex agents, and they are also faced with the exploration-diversity conflict.
+To this end, our ComSD proposes to design strong incentives through contrastive learning and presents a novel dynamic weighting algorithm (SMW) to overcome the exploration drop brought by explicit conditioned entropy maximization (i.e., diversity reward).
 CIC
-(Laskin et al., 2022b) is a state-of-the-art robot behavior discovery method.
-It first introduces contrastive learning (Chen et al., 2020) into unsupervised skill discovery.
+Laskin et al.
+[2022b] is a state-of-the-art robot behavior discovery method.
+It first introduces contrastive learning Chen et al.
+[2020] into unsupervised skill discovery.
 It chooses the second MI decomposition, Eq.
@@
 The encoder learned by contrastive learning is further used for APT reward improvement.
-The behaviors produced by CIC are of high activity but not distinguishable.
-Different from CIC, we employ the contrastive results as diversity intrinsic rewards for explicit conditioned entropy maximization to improve behavioral diversity, with SMW to balance two entropy estimations for exploratory ability maintenance.
+The behaviors produced by CIC are of high activity but low diversity.
+Different from CIC, we employ the contrastive results as a novel diversity reward for explicit conditioned entropy maximization to improve behavioral diversity, with SMW for exploration maintenance.
 BeCL
-(Yang et al., 2023) is another state-of-the-art method on URLB (Laskin et al., 2021) and 2D exploration (Campos et al., 2020).
+Yang et al.
+[2023] is another state-of-the-art method on URLB Laskin et al.
+[2021] and maze exploration Campos et al.
+[2020].
 It tries to mitigate the exploitation problem in CIC by a novel MI objective, $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill.
@@
 Appendix D Additional Analysis
-D.1 What Skills Do ComSD and Competitive Baselines Discover? & Why does ComSD exhibit better adaptation performance than others?
+D.1 Additional Diversity Analysis
+In this section, we further analyze the behavioral diversity of each method.
+In Section 4.3, we use Mean State (MS) to represent one skill, and calculate the MS-coverage over several skills (selected by grid sweep Laskin et al.
+[2022b]) to represent the behavioral diversity of one method.
+In addition to MS-coverage, we further define another two metrics for behavioral diversity evaluation: MS-coverage-aknn and MS-range.
+Here we provide the concrete computation of these three metrics.
+MS-coverage: For one method, we calculate the MS Euclidean distance between each skill and its k-th-nearest neighbor skill, then take the mean.
+MS-coverage-aknn: For one method, we calculate the mean MS Euclidean distances between each skill and its all k-nearest neighbor skills, then further take the mean.
+MS-range: For one method, we report the maximum MS range of selected skills.
+The comparison between all 6 methods on behavioral diversity is shown in Figure 7.
+ComSD has huge advantages on two coverage metrics, demonstrating that its behaviors have the largest coverage in the potential skill space.
+Perhaps the most interesting phenomenon is that ComSD is also the most competitive method on MS-range.
+In fact, it’s hard to represent a dynamic behavior by MS (two different dynamic behaviors may have the same mean state), which puts highly exploratory methods (ComSD and CIC) at a disadvantage on these three metrics.
+In this case, CIC is at the bottom of all three metrics, while ComSD still obtains state-of-the-art evaluation results.
+This effectively demonstrates that ComSD produces behaviors of much higher diversity than all other baselines.
+Figure 7: The comparison between all 6 methods on 3 diversity metrics. ComSD discovers behaviors of much more diversity than other baselines.
+D.2 Skill Visualization: What Skills Do ComSD and Competitive Baselines Discover?
 Figure 8: Visualization for representative behaviors discovered by CIC and APS.
-AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4.
-Skill AKD ranges from 7 to 10 for CIC and 0 to 2 for APS, which means they can’t discover qualified behaviors at different activity levels.
+AKD is a state entropy estimator used to evaluate skill exploration (activity), which we define in Section 4.3.
+Skill AKD ranges from 7 to 10 in CIC and 0 to 2 in APS, which means they can’t generate behaviors at different exploratory levels.
+In addition, CIC’s behaviors are also of low diversity.
 Figure 9: Visualization for representative behaviors discovered by ComSD.
-AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4.
-Skill AKD ranges from 0 to 10 for ComSD.
-The results demonstrate that our ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot.
-We provide the skill visualization and corresponding skill AKD of the two most competitive baselines, APS (Liu & Abbeel, 2021a) and CIC (Laskin et al., 2022b), in Figure 6.
+AKD is a state entropy estimator used to evaluate skill exploration (activity), which we define in Section 4.3.
+Skill AKD ranges from 0 to 10 in ComSD.
+The results verify that our ComSD can produce diverse behaviors at different exploratory levels, which recent advanced methods (CIC and APS) cannot.
+We provide the skill visualization and corresponding skill exploration (AKD) of the two most competitive baselines, APS Liu and Abbeel [2021a] and CIC Laskin et al.
+[2022b], in Figure 8.
 The visualization and skill AKD of our ComSD are shown in Figure 9.
-AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4.
-CIC is able to produce continuous movements of high activity, but it can’t generate behaviors at other activity levels (AKD range of CIC’s skills is 7-10).
-In addition, CIC suffers from insufficient exploitation, i.e., the generated skills are indistinguishable and homogeneous.
+AKD is used to represent skill exploration (activity), which we define in Section 4.3.
+CIC is able to produce continuous movements of high exploration, but it can’t generate behaviors at other activity levels (AKD range of CIC’s skills is 7-10).
+In addition, CIC suffers from insufficient diversity, i.e., the generated skills are indistinguishable and homogeneous.
 CIC’s skills all tend to achieve dynamic flipping, which is consistent with the good initial score on walker flip in skill combination (see Section 4.2).
-However, the behavioral indistinguishability makes it difficult for meta-controllers to learn ideal combinations for a competitive final score.
-APS can generate diverse behaviors, but it suffers from lazy exploration and also can’t generate behaviors at different activity levels (AKD range of APS’s skills is 0-2).
-In general, poor exploration causes poor performance in skill finetuning evaluation.
-In skill combination, the diversity allows the meta-controller to complete downstream tasks through the combination of unqualified learned skills, which coincides with the upward trend of APS training curves (see Section 4.2).
-In summary, previous advanced methods can’t provide a good balance between behavioral quality and diversity, thus failing to exhibit competitive results across different downstream adaptation evaluations.
-By contrast, our ComSD can produce diverse behaviors at different activity levels (AKD range of ComSD’s skills is 0-10), including flipping, lying down, struggling at different speeds, various postures, and so on.
-This explains why our ComSD can achieve state-of-the-art adaptation performance across both kinds of downstream tasks while other methods cannot.
-D.2 Behavioral Diversity Analysis
-In this section, we try to analyze the behavioral diversity of each method.
-For a skill, we use it to sample 1k states and calculate the Mean State (MS) over all sampled states.
-MS can partially represent the skill to some extent.
-For each method, we uniformly sample 41 different walker skills.
-Over 41 skills, we compute the K-Th-nearest-neighbor MS Distance (KTMSD) and the All-K-nearset-neighbor MS Distance (AKMSD) for skill entropy estimation.
-KTMSD computes the k-th-nearest-neighbor MS Euclidean distance of each skill and takes the mean, while AKMSD considers all the k-nearest neighbors of each skill and takes the mean distance of MS.
-(Note that in Section 4.4&Appendix D.1, KTD and AKD are employed for state entropy estimation of one skill, while KTMSD and AKMSD are used for skill entropy estimation of one method in this section.) KTMSD and AKMSD can partially evaluate the skill coverage of one method.
-In addition, we also calculate the MS Range (MSR) for each method.
-These metrics are all related to behavioral diversity.
-The comparison between all 6 methods on behavioral diversity is shown in Figure 8, demonstrating that ComSD has huge advantages on skill entropy (KTMSD and AKMSD).
-ComSD is also the most competitive method on MSR.
-In fact, it’s hard to represent an exploratory skill by only MS (exploratory behaviors have much more visited states than static postures), which puts highly exploratory methods (ComSD and CIC) at a disadvantage on these 3 metrics.
-In this case, ComSD still obtains state-of-the-art evaluation results, which demonstrates that the behavior set discovered by ComSD is of much higher diversity and coverage than other baselines.
-Figure 8: The comparison between all 6 methods on behavioral diversity. ComSD discovers a much more diverse skill set than other baselines.
+However, the low behavioral diversity makes it difficult for meta-controllers to learn ideal combinations for a competitive final score.
+APS can generate diverse robot skills, but it suffers from lazy exploration and also can’t generate behaviors at different activity levels (AKD range of APS’s skills is 0-2).
+Compared with CIC, the higher diversity allows the meta-controller to partially complete downstream tasks, which coincides with the upward trend of APS training curves (see Section 4.2).
+However, the poor exploration causes poor performance of APS in skill finetuning, simultaneously.
+In summary, previous advanced methods can’t provide a good balance between behavioral exploration and diversity, thus failing to exhibit competitive results across different downstream adaptation evaluations.
+By contrast, our ComSD can produce diverse behaviors at different exploratory levels (AKD range of ComSD’s skills is 0-10), including flipping, lying down, struggling at different speeds, various postures, and so on.
+ComSD’s huge superiority in behavioral quality explains why it achieves state-of-the-art adaptation performance across both kinds of downstream tasks, while other methods cannot.